A Bibliography of Publications in Open Systems Today Magazine

Nelson H. F. Beebe Department of Mathematics

University of Utah Salt Lake City, UT 84112

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 September 2010 Version 1.03

Title word cross-reference

#10 [Ano94-189]. #11 [Ano94-190]. #9 [Ano94-194].

11-part [Ano94-191].

2 [Ano95-30]. 2-D [Ano95-30]. 2.0 [Ano94-60]. 25 [Ano94-209]. 25-Mbps [Ano94-209]. 28.8 [Ano95c]. 28.8-Kbps [Ano95c]. 295 [Ano94-67].

3 [Ano94-182]. 3-D [Ano94-182]. 32-bit [Ano94-158]. 35om [Ano94a]. 3Com [Ano94b]. 3CT [Ano95-29].

4.1 [Ano94-154]. 400s [Ano94-66]. 486-based [Ano95t]. 4GL [Ano94-220]. 4GLs [Ano95-46].

500 [Ano94-38]. 500-Kbps [Ano94-38].

6000s [Ano94-84]. 64-bit [Ano94-184].

75 [Ano94g]. 75-MHz [Ano94g].

95 [Ano94-213].

abilities [Ano94-152]. able [Ano94-234]. aboard [Ano94-214]. access [Ano94-150, Ano94-195]. accessing [Ano95-52]. according [Ano94-53]. accounting [Ano94-56]. accounts [Ano94-107]. Acrobat [Ano94t]. across [Ano94-220]. ad [Ano94x]. adapter [Ano94a]. adapters [Ano95o]. add [Ano94-234]. adding [Ano94-74]. addition [Ano95q]. additions [Ano95a]. address [Ano94-62]. addresses [Ano94-198]. adds [Ano94-115]. adhere [Ano94-83]. Admin [Ano94-198]. Administration [Hun94a]. administrative [Ano95f]. administrator [Ano95-60]. Adobe [Ano94t]. advantages [Ano94-232]. advertisement [Ano94x]. advertising [Ano95-45]. advice [Ano95-45]. Advisor [Lyt94a]. advocate [Ano94-176]. against [Ano94-203]. agent [Ano94-112]. agent-based [Ano94-112]. aggressively [Ano94-234]. aim [Ano94-225]. aimed [Ano94-146]. AIX [Ano94-73]. alerts [Ano94-30]. allocated [Ano94-198]. allow [Ano95l]. allowing [Ano94-221]. allows [Ano94-74]. allure [Ano94u]. alpha [Ano94-160]. ALR [Ano94-148]. also [Ano94u]. alternative [Ano94-213]. alternatives [Ano94-36]. Although [Ano94u]. among [Ano94-179, Ano95l]. amount [Ano94-114]. analysis [Ano94-178]. analytical [Ano95d]. analyzing [Ano95-49]. announce [Ano94-77]. announced [Ano94-73]. announcement [Ano94-234]. announcements [Ano95l]. announcing [Ano94-164]. ANSI [Ano95-61]. ANSI- [Ano95-61]. antique [Ano94-109]. anymore [Ano94-43]. AnyNet [Ano94-74]. API [Ano95-50]. APIs [Ano94-101]. aplomb [Ano94-47]. app [Ano94-183]. apparently [Ano94-48]. appears [Ano94-106]. applaud [Ano94-29]. Apple [Ano94-168, Ano94y]. application [Ano94-234, Ano95-66]. application-suite [Ano94-234]. applications [Ano94-168, Ano94-35]. approach [Ano94-229]. apps [Ano94-234, Ano94-228]. April [Ano95m]. Architecture [Ano94-138, Ano94-140]. archiving [Ano94-43, Ano94-223]. area [Ano94-192]. array [Ano95g]. AS/ [Ano94-66]. ASIC [Ano94b]. ASIC-based [Ano94b]. AST [Ano94-144]. Athena [Ano94-47]. ATM [Ano94-33]. attendees [Ano94-168]. audio [Ano94-71]. automate [Ano95f]. automated [Ano95-43]. available [Ano94-157, Ano95-61]. Avalon [Ano94-162]. AViiON [Ano94-42]. Avistar [Ano94k]. awaited [Ano94t]. awaits [Ano94-189]. AXP [Ano94c].

back [Ano95-54, Ano94-87].

backbone [Ano94-119]. backed [Ano94-33]. bag [Ano94-132]. bandwidth [Ano94-192]. Bank [Ano94-33]. banker [Ano94-68]. Banyan [Ano94-28]. barristers [Ano94-205]. base [Ano94o]. based [Ano94-168, Ano94-86, Ano94-172]. basing [Ano95w]. Bay [Ano94-31]. BBN [Ano95k]. be [Ano94-234]. Bear [Ano94-68]. beat [Ano94-173]. because [Ano94-224]. been [Ano94-168]. before [Ano94-199]. began [Ano95-68]. begins [Ano95b]. being [Ano94-114]. benchmark [Ano94c]. benchmarks [Ano95-64]. benefits [Ano94m]. best [Ano94-166]. beta [Ano94-207, Ano94-29]. better [Ano94-53]. between [Ano94-111]. Beverages [Ano95s]. beyond [Ano94-140]. bid [Ano94-176]. big [Ano94-49]. billing [Ano94-78]. bit [Ano94-184]. Blue [Ano95q, Ano95q]. Blyth [Ano94-174]. Bob [Ano94-32]. body [Ano94x]. book [Ano95-45]. boost [Ano94-169]. Borland [Ano94-52]. both [Ano94-228, Ano94-232]. bottlenecks [Ano94-188]. box [Ano94-109, Ano94-61]. brain [Ano95-67]. brain-weary [Ano95-67]. branding [Ano94-221]. break [Ano95-67]. briefing [Ano94-153]. bring [Ano94-215]. bringing [Ano94s]. brings [Ano95h]. British [Ano94x]. broadly [Ano94-83]. Broker [Ano94-138]. browser [Ano94-36]. bugs [Ano94-216]. build [Ano94-57]. Builder [Ano94-101]. builders [Ano94-157]. building [Ano94-174]. builds [Ano94-170]. built [Ano94-37]. built-in [Ano94-37]. Bull [Ano94-63]. bundling [Ano94-167]. Burlington [Ano94w]. business [Ano94-178, Ano94-30]. buying [Ano94-224].

C [Ano94-85]. Cabletron [Ano95r]. cache [Ano95n]. Cadbury [Ano95s]. call [Ano94-107]. Calluna [Ano94-169]. came [Ano94-153]. can [Ano94-201]. Canon [Ano95t]. capabilities [Ano94-70, Ano94-37]. capability [Ano94-58]. capacity [Ano94-169]. card [Ano94-168, Ano94j]. cards [Ano94-197]. catch [Ano94-31]. cc [Ano94-116]. CD [Ano94-42]. CDPD [Ano94-179]. CDs [Ano94-223]. cellular [Ano94-179, Ano94-31]. center [Ano95a]. CEO [Ano94-32]. challenge [Ano94z]. challenges [Ano95e]. change [Ano95-64]. characters [Ano94-116]. Chase [Ano94-33]. chassis [Ano94a]. checkbooks [Ano94-168]. chip [Ano94-217, Ano94-233]. chips [Ano94-184]. choice [Ano94-208]. choose [Ano94-108]. chose [Ano94-55]. Cisco [Ano94-34]. citing [Ano94-29]. Clarify [Ano94-35]. clearly [Ano95b]. click [Ano94-102]. client [Ano94-101]. client-server [Ano94-187]. clients [Ano95-69, Ano95-69]. clock [Ano94-69]. closely [Ano95-62]. closer [Ano94-27]. coalition [Ano94v]. Coat [Ano94w]. COBOL [Ano95-46]. code [Ano94-201]. collaborative [Ano94-80]. Collabra [Ano94-28, Ano95u]. collectively [Ano95-60]. color [Ano94-67]. combine [Ano94-222]. come [Ano94y]. comes [Ano94l]. comet [Ano95-68]. commands [Ano95r]. Commerce [Ano94-189]. commercial [Ano94-146, Ano94-36]. Common [Ano94-138]. Communication [Ano94-139]. communications [Ano94-117]. companie [Ano94-137]. companies [Ano94-137]. company [Ano94-170]. Compaq [Ano94-148, Ano94-37]. Comparative [Ano95v]. compares [Ano94-166]. comparison [Ano94o]. compatibility [Ano94y]. compatible [Ano94-168]. compete [Ano94-36]. competition [Ano95-47, Ano94c]. competitive [Ano94-43]. complex [Ano94-166]. compound [Ano94-134]. comprehensive [Ano94-174]. computer [Ano94-137]. computers [Ano94-40, Ano94g]. Computing [Ano94r, Ano94-50]. concern [Ano94-105]. concerns [Ano94-212]. concludes [Ano94-190]. conference [Ano94-105, Ano95-67]. conference-goers [Ano95-67]. conferencing [Ano94-104]. configuration [Ano94-114]. configured [Ano95r]. connect [Ano94i]. connection [Ano94m]. connections [Ano94-175]. connectivity [Ano94-101, Ano94-38]. considering [Ano94o]. considers [Ano94-191]. consist [Ano94-156]. console [Ano94-207]. consoles [Ano94-112]. consulting [Ano94-140]. Control [Ano94-216]. converts [Ano95-31]. cool [Ano94g]. cooler [Ano94-233]. COPD [Ano94-39]. coprocessor [Ano94j]. CORBA [Ano94-60]. core [Ano94-32]. Cornerstone [Ano95k]. cornerstones [Ano94-158]. corporate [Ano94-86]. cost [Ano94-192]. counters [Ano94-66]. country [Ano94p]. country-and-western [Ano94p]. course [Ano94-230]. cozy [Ano94-108]. crafts [Ano94-109]. created [Ano95-66]. creating [Ano94u]. cross [Ano94-228]. cross-platform [Ano94-228]. Cruz [Ano94-215]. custom [Ano94q]. customer [Ano94-35]. customer-service [Ano95q]. customers [Ano94-62]. customize [Ano94-167]. cut [Ano94-226]. cuts [Ano94-66]. Cyberleaf [Ano95-31].

D [Ano94-182]. dams [Ano94-230]. Data [Ano94-179, Ano94-42]. data-analysis [Ano95k]. data-staging [Ano94-171]. database [Ano94-166, Ano94-170]. databases [Ano94-111, Ano94-86]. datacom [Ano94-39]. Dawn [Ano94-44]. days [Ano94-137]. DB2 [Ano94-77]. dBase [Ano94-52]. DBMS [Ano94-64]. DBs [Ano94-167]. DCA [Ano94-45]. DCE [Ano94-103]. deadbeat [Ano94-107]. debate [Ano94-232]. debuts [Ano94-46]. DEC [Ano94c, Ano94-46]. DECathena [Ano94-47]. decision [Ano94z]. decisions [Ano94-224]. DECpc [Ano94-208]. Defect [Ano94-216]. defensive [Ano94z]. define [Ano95-46]. deliver [Ano94-79]. Dell [Ano94-148]. demonstration [Ano94-168]. Department [Ano94-55]. depend [Ano94-217]. deployment [Ano95-51]. depth [Ano94-229, Ano95-32]. design [Ano94-155]. designed [Ano94-145]. designs [Ano94-184]. desktop [Ano94-151, Ano94-144]. desktops [Ano94-149]. desperate [Ano94-107]. despite [Ano94-52]. destinations [Ano94-193]. developed [Ano94-114]. developers [Ano94-228]. developing [Ano95t]. development [Ano94-102]. development-tool [Ano94-65]. device [Ano94-38]. devices [Ano94-76]. DHCP [Ano94-198]. diagnosis [Ano94-45]. dial [Ano94-186]. dial-up [Ano94-186]. did [Ano94-153]. different [Ano94-225]. Digital [Ano94-179, Ano94-48]. Digitalk [Ano95v]. directory [Ano94-119]. disappear [Ano95-64]. discontinue [Ano94-183]. disk [Ano94-169]. disk-drive [Ano94-169]. DiskShare [Ano94-175]. disparate [Ano94-111]. distribute [Ano94-220]. Distributed [Ano94r, Ano94-50]. distribution [Ano94-45]. diverse [Ano94-136]. divining [Ano94-137]. DMBS [Ano94-49]. DMI [Ano94-105]. do [Ano94-191]. document [Ano94t]. documents [Ano95-31]. Documentum [Ano94-51]. doesn’t [Ano94k]. dog [Ano94-106]. doing [Ano95-45]. don’t [Ano94-40]. DOS- [Ano94-168]. Dow [Ano94-170]. down [Ano94-177]. drive [Ano94-169]. drives [Ano95n]. dropping [Ano94-134]. dual [Ano94-149]. dual-processor [Ano94-149]. dubbed [Ano94-78]. due [Ano94-81]. duty [Ano94-145]. dynamic [Ano94-114]. dynamically [Ano94-198].

E-mail [Ano94-180, Ano95-48].

each [Ano94-35].

early [Ano94-161, Ano94-52].

eases [Ano95-61]. easier [Ano94-113].

easily [Ano95-48, Ano94-167]. easy [Ano94-195, Ano94-40]. easy-to-build [Ano95-44]. eclipse [Ano94-211]. EDA [Ano94-101]. EDA/ [Ano94-101]. Edge [Ano94-216]. edged [Ano94c]. editor [Ano95-54]. education [Ano95d]. effective [Ano94-192]. efficiently [Ano94-111]. eight [Ano94v]. eight-member [Ano94v]. electronic [Ano94-180]. Electronics [Ano95w]. eliminate [Ano94-188]. embrace [Ano94-202]. emerging [Ano94-198]. empower [Ano94-205]. emulation [Ano94-151]. EMX [Ano94-119]. Encore [Ano95g]. encounters [Ano94-87]. encouraging [Ano94-146]. encryption [Ano95-53]. end [Ano94-27]. ends [Ano94-180]. Energy [Ano94-55]. engine [Ano94-117, Ano94-135]. engineers’ [Ano94l]. engineers’-get-together [Ano94l]. enhanced [Ano94-101]. enhancements [Ano94-215]. enhances [Ano94q]. enough [Ano94g]. enter [Ano94-207]. enterprise [Ano94-57, Ano95-62]. enterprise-wide [Ano94-57]. entice [Ano94-66]. entry [Ano94-208, Ano94-53]. entry-level [Ano94-208, Ano94-53]. environment [Ano94-216, Ano94-234]. environments [Ano94-226]. Equipment [Ano94-48]. equivalents [Ano95-31]. era [Ano94-44]. Escala [Ano94-63]. Eskimo [Ano94-218]. especially [Ano94-224]. Ethernet [Ano94-197]. even [Ano94-172]. Ever [Ano94-54]. examined [Ano94-229]. Exchange [Ano94-135]. exhibited [Ano94-64]. existing [Ano94-83]. expands [Ano94-73]. expected [Ano94-209]. experience [Ano94e]. experts [Ano94-190]. Explorations [Bea94a]. Expo [Ano94-85, Ano94-181]. eye [Ano95-68]. Eye2eye [Ano95-57].

F [Ano94-114]. F-22 [Ano94-114].

Faced [Ano94-55]. Factory [Ano94w]. failover [Ano94-58]. fall [Ano94-103]. far [Ano94-222]. fast [Ano94-224]. faster [Ano95n]. fault [Ano94-69, Ano95x]. fault-resilient [Ano94-69, Ano95x]. faxing [Ano94-68]. FDDI [Ano94-58]. feature [Ano94-144]. featured [Ano95-58]. features [Ano94-101, Ano94y]. few [Ano94-233]. Fidelity [Ano94-57]. field [Ano95-32]. fighter [Ano94-114]. financial [Ano94-162, Ano94-57]. find [Ano94-62]. finding [Ano94-235]. Firefox [Ano94-58]. FirePower [Ano94-149]. firewall [Ano94d]. firm [Ano94-205]. first [Ano94-81]. five [Ano95-54]. flexibility [Ano94-196]. floating [Ano95-29]. floating-point [Ano95-29]. flood [Ano94-172]. flurry [Ano95l]. focus [Ano94n, Ano94-46]. following [Ano95-62]. form [Ano94-160]. formed [Ano94-59]. forms [Ano94-231]. forthcoming [Ano94-80]. Forum [Ano94-179]. founded [Ano94-179]. four [Ano94-111, Ano94-60]. four-module [Ano94-111]. framework [Ano94-65]. Francisco [Ano94-31]. Frankenberg [Ano94-32]. free [Ano95-52]. freedom [Ano94-108]. freely [Ano94-217]. freeware [Ano94-139, Ano94-36]. French [Ano94-87]. fret [Ano94-54]. friends [Ano94-54]. frills [Ano95c]. front [Ano94-234]. ftp [Ano95p]. Fujitsu [Ano94-61]. full [Ano94m]. functions [Ano94-222]. fundamental [Ano95b]. future [Ano94-212, Ano94-190].

gain [Ano94-43]. gaining [Ano95-62].

Galaxy [Ano94-228].

gateway [Ano94-58, Ano94-55].

GE [Ano94-62]. General [Ano94-42]. generation [Ano94-76]. get [Ano94l]. gets [Ano94-107]. giant [Ano94-57]. GIS [Ano94-37]. give [Ano94-52]. gives [Ano94h]. giving [Ano94m]. Global [Ano95c]. go [Ano94-106]. God [Ano94-153]. goers [Ano95-67]. goes [Ano94-212]. going [Ano95-53]. Good [Ano94-176]. got [Ano94-68]. government [Ano94-176]. Graphics [Ano94-182]. great [Ano94-82]. groped [Ano94-168]. Group [Ano94-214]. group-conferencing [Ano95u]. Groupe [Ano94-63]. groups [Ano95-60]. growing [Ano94-37]. grows [Ano94-200, Ano95q]. GUI [Ano94-174]. guide [Ano94-157]. GUIs [Ano95-46].

hair [Ano94-177]. handle [Ano94-68]. happen [Ano94-165]. happens [Ano94-193]. hard [Ano94-235]. hardware [Ano94-217, Ano94-208]. harmony [Ano94-132]. having [Ano94-105]. head [Ano94-176, Ano95v]. head-to-head [Ano95v]. held [Ano94-153]. help [Ano94d]. here [Ano94-213, Ano94-64]. heterogeneous [Ano94-50]. Hewlett [Ano94-65]. high [Ano94-52]. high-performance [Ano94-72]. high-tech [Ano94e]. hold [Ano94-224]. home [Ano94-62]. Honeywell [Ano94d]. HP [Ano94-66, Ano94-70]. HP-UX [Ano95y]. HTML [Ano95-31]. hub [Ano94a]. Hubble [Ano95-68]. hubs [Ano94-211]. huge [Ano94-47]. Human [Ano94-71]. hydroelectric [Ano94-230]. hyped [Ano94-31].

I/O [Ano95y]. IBM [Ano94-160, Ano94-74]. Icon [Ano94-86]. Icon-based [Ano94-86]. ideal [Ano94-222]. identify [Ano94d]. Illustrator [Ano94s]. Ilog [Ano95-30]. impact [Ano94-231]. implementation [Ano94-209, Ano94-29]. implementations [Ano94-60]. important [Ano94-201]. Improv [Ano94-183]. improved [Ano94-115]. in-depth [Ano94-229, Ano95-32]. include [Ano94-135]. includes [Ano94-198]. including [Ano94-164]. incorporate [Ano94-71]. incorporated [Ano94-185]. incorporation [Ano95m]. increased [Ano95-69]. increases [Ano95-47]. increasing [Ano94-146]. independence [Ano94-103]. independent [Ano94-48]. indexing [Ano94-203]. industry [Ano94-63, Ano94-88]. Information [Ano94-150, Ano94-101]. Informix [Ano94-102]. InForms [Ano95-48]. infrastructure [Ano94v]. initiative [Ano94-214]. inroads [Ano94-133]. Institute [Ano94-178]. Instruments [Ano95w]. integraph [Ano94-175]. integrate [Ano94-194]. integrated [Ano94-210]. integrates [Ano94-61]. integrating [Ano94-59]. integration [Ano94t]. Intel [Ano95-47, Ano94-104]. Interactive [Ano94-197]. interchangeable [Ano95-51]. interface [Ano94h, Ano94-51]. interfere [Ano94k]. Intergraph [Ano94-144]. Interleaf [Ano95-31]. Internet [Ano94-135, Ano94w]. Interop [Ano94-164]. interoperability [Ano94-179, Ano94-151]. interoperate [Ano95-66]. InterPPP [Ano94-219]. introduce [Ano94-181]. introduced [Ano94-111]. introduces [Ano94-65]. introducing [Ano94-140]. intruders [Ano94d]. investigate [Ano95x]. investment [Ano94-68]. Investments [Ano94-57]. IP [Ano94h]. ISDN [Ano94-115]. isn’t [Ano94-43]. issues [Ano94-189]. itself [Ano94-235]. IXI [Ano95-57].

JAM [Ano94f]. jazzing [Ano94-131]. job [Ano94-166]. join [Ano94-37]. jousting [Ano94-156]. juke [Ano94-42]. juke-box [Ano94-42]. Jupiter [Ano95-68]. just [Ano94-201]. JYACC [Ano94f].

Karl [Ano94-54, Ano94-109]. Kbps [Ano94-38]. keep [Ano94-131]. kernel [Ano95y]. key [Ano94-41]. kind [Ano95-67]. know [Ano94-40].

Labs [Ano94-231]. lack [Ano94-55]. Lakes [Ano94-80]. LAN [Ano94-191, Ano94-110]. lances [Ano94-156]. lands [Ano94p]. languages [Ano94-210]. LANPlex [Ano95a]. LANs [Ano95c]. laptop [Ano95h]. larger [Ano95n]. laser [Ano94-67]. last [Ano95-68]. late [Ano94-54]. latest [Ano94c]. law [Ano94-205]. leaders [Ano94-59]. leading [Ano94-202]. leading-edge [Ano94-202]. leads [Ano94-85]. least [Ano94-33]. legacy [Ano94-147]. Legent [Ano95f, Ano94-111]. less [Ano94-69]. let [Ano94-180]. lets [Ano94-226]. letting [Ano94-116]. level [Ano94-208, Ano94-53]. leverages [Ano94-107]. Liant [Ano94-113]. license [Ano94n]. licensed [Ano94-217]. lighter [Ano95e]. like [Ano94-49]. line [Ano94-101, Ano95c]. lines [Ano94-182]. link [Ano94-60, Ano94-38]. linking [Ano94-147]. links [Ano95-57]. Lisp [Ano94-210]. little [Ano94z].

Live [Ano95-37]. loaded [Ano94-53]. local [Ano94-167]. locker [Ano94-218]. Lockheed [Ano94-114]. Locus [Ano94-226]. lode [Ano94-215]. logic [Ano95-46]. long [Ano94t, Ano94-116]. long-awaited [Ano94t]. look [Ano94-192]. looked [Ano94-33]. looking [Ano94-222, Ano94-115]. looks [Ano94-231]. lot [Ano94-69]. lots [Ano94-64]. Lotu [Ano94-118]. Lotus [Ano94-183, Ano94-116]. love [Ano94-54]. low [Ano94-192]. low-cost [Ano94-192]. low-end [Ano94-75]. low-voltage [Ano94g]. lower [Ano95t]. lower-cost [Ano95t].

Mac [Ano94-168]. machines [Ano94-144]. Macintosh [Ano94-219]. MacPPP [Ano94-219]. Macs [Ano94-224]. made [Ano94-202]. mail [Ano94-135, Ano94-120]. mainframe [Ano94c, Ano94-200]. mainframe-to-LAN [Ano94-200]. mainframes [Ano94-194, Ano94-76]. maintains [Ano94x]. major [Ano94-137, Ano94-131]. make [Ano94-187]. Maker [Ano95-50]. makes [Ano94-175]. making [Ano94-133, Ano94-132]. man [Ano94p]. manage [Ano94-112]. management [Ano94-222]. management-system [Ano94-51]. managers [Ano94-194]. manages [Ano94-216]. managing [Ano94-170]. Manhattan [Ano94-33]. manipulation [Ano94-101]. manuals [Ano94-54]. manufacturing [Ano94-162]. Many [Ano94-233]. MAPI [Ano94-117]. mapping [Ano95-31]. Marion [Ano94-170]. market [Ano94-156]. market-leading [Ano94-67]. marketing [Ano94-66]. marketplace [Ano95-47]. marks [Ano94-52]. MasPar [Ano94-72]. Massively [Ano94-133]. master [Ano94n]. matching [Ano94-66]. materializing [Ano94-224]. matures [Ano94-196]. may [Ano94z]. Mbps [Ano94-209]. means [Ano94-137]. meat [Ano94-218]. meet [Ano94-138]. Meeting [Ano95-50]. melded [Ano95s]. member [Ano94v]. memory [Ano95n]. Mercury [Ano95-43, Ano95-43]. Merge [Ano94-226]. Merrell [Ano94-170]. message [Ano94-193]. messages [Ano94-116]. messaging [Ano94-135, Ano94-232]. metering [Ano94-45]. MHz [Ano94g]. microkernels [Ano94-158]. Microsoft [Ano94-106, Ano94-135]. middle [Ano94-70]. middleware [Ano94-101, Ano94-136]. midrange [Ano94b, Ano94-110]. migrate [Ano94-66]. Mike [Azz94a]. MIME [Ano95m]. MiniStor [Ano94-169]. misleading [Ano94x]. mobile [Ano94h]. Model [Ano95-29, Ano94-159]. models [Ano94-75, Ano94-83]. modem [Ano95c]. modular [Ano95q]. modularity [Ano94-158]. module [Ano94-111]. modules [Ano94b]. money [Ano94-69]. monitor [Ano94-227]. month [Ano94-181, Ano94-73]. Mosaic [Ano94n, Ano94-139]. most [Ano94-187]. mostly [Ano94-156]. mother [Ano94-215]. move [Ano94z]. moved [Ano94-234]. mover [Ano94-111]. moving [Ano94-27, Ano94-140]. much [Ano94u]. multidatabase [Ano95-51]. multilevel [Ano94-30]. multimedia [Ano94-135, Ano94-131]. multiplatform [Ano94-35]. multiple [Ano94-228]. multiple- [Ano94-228]. multiply [Ano94-192]. multiprocessing [Ano94-81]. multiprocessor [Ano94a, Ano94-141]. multitasking [Ano94-144]. multithreaded [Ano94-141]. must [Ano94-136]. MVS [Ano94-85]. My [Azz94e].

name [Ano95-60, Ano94-221].

names [Ano94-201]. native [Ano94-135]. NCSA [Ano94-36, Ano94n]. NDS [Ano94-29]. nearly [Ano94-69]. NEC [Ano94-144]. need [Ano94-199]. needs [Ano94o]. neighbor [Ano94-109]. net [Ano94-165, Ano94-142]. NetBIOS [Ano94-74]. NetManaged [Ano94-115]. NeTpower [Ano94-149]. nets [Ano94i, Ano94-194]. Netscape [Ano94-212]. NetView [Ano94-73]. NetViews [Ano94-143]. NetWare [Ano94-197, Ano94-48]. network [Ano94-191, Ano94-104]. network-in-a-box [Ano94-109]. networking [Ano94y, Gas94a]. networks [Ano94-74, Ano94-34]. NetWorld [Ano94-164]. NetWorld-Interop [Ano94-164]. NewEra [Ano94-102]. newly [Ano95y]. news [Ano94-85]. next [Ano94l]. next-generation [Ano94l]. NextStep [Ano95t]. NFS [Ano94-61]. no [Ano94-168]. no-frills [Ano95c]. non [Ano95-53]. non-Oracle [Ano95-53]. Northern [Ano94-151]. Note [Ano94-152]. notebook [Ano94g]. notebooks [Ano94-233]. Notes [Ano94-150]. Novell [Ano94-32, Ano94-153]. November [Ano94s]. NOV’IX [Ano94-58]. NT [Ano94-148, Ano95l]. NT-NFS [Ano94-175]. number [Ano94-37]. numbers [Ano95-29].

object [Ano94-102, Ano94-157].

object-based [Ano95-44]. object-oriented [Ano94-102, Ano94-157]. Object.station [Ano95t]. Objectivity [Ano95-66]. Objectivity/ DB [Ano95-66]. objects [Ano94u, Ano94-158]. obtains [Ano94n]. ODBC [Ano94-56]. Odyssey [Atk94g]. off [Ano94-33]. off-line [Ano94-30]. offenses [Ano94x]. offer [Ano94-104]. offered [Ano95-37]. offering [Ano94-138]. offers [Ano94r]. office [Ano94-224]. often [Ano94-132]. OLE [Ano94o, Ano94-134]. Omnis [Ano94-174]. on-line [Ano94-27, Ano95-49]. once [Ano94-31]. once-hyped [Ano94-31]. one [Ano94-222]. one-man [Ano94p]. only [Ano94-168]. Ontos [Ano94-159]. OOPSLA [Ano94-155]. Open [Lyt94a]. OpenDoc [Ano94o, Ano94-160]. OpenLink [Ano95-51]. openness [Ano94-82]. opens [Ano94-150]. OpenStep [Ano94-161]. OpenView [Ano94-70]. operating [Ano94-197]. Operation [Ano94-215, Ano94-30]. operations [Ano95f]. optimized [Ano95-43]. options [Ano94-101]. Oracle [Ano94-49, Ano94-164]. Oracle7 [Ano95h]. Ordering [Ano94-165]. Oregon [Ano95q]. organizational [Ano94-46]. organizations [Ano94-136]. oriented [Ano94-102, Ano94-157]. OS/ [Ano94-160]. OSecs [Ano95-51]. OSes [Ano94-158]. OST [Ano94-192, Ano94-166]. OSTs [Ano94-191]. other [Ano94-52]. our [Ano94-166, Ano95-54]. outlines [Ano94-154]. overhaul [Ano95-64]. overhauled [Ano94-101]. own [Ano94-80].

Pacer [Ano95-55]. package [Ano94h, Ano94f]. packaged [Ano94-140]. packages [Ano94-219, Ano94-231]. Packard [Ano94-65]. Packet [Ano94-179]. pager [Ano94-116]. pages [Ano94-62]. paperless [Ano94-231]. parallel [Ano94-133]. ParcPlace [Ano95-66]. part [Ano94-191]. partition [Ano94-220]. partnerships [Ano94-137]. parts [Ano94-220]. party [Ano94-234, Ano95-63]. passes [Ano94-111]. past [Ano95-54]. paste [Ano94-226]. paths [Ano95y]. PC [Ano94y, Ano94-216]. PC-Admin [Ano94-198]. PC-based [Ano95-46]. PC-NFS [Ano94-175]. PCExpo [Ano94-168]. PCMCIA [Ano94-197, Ano94-169]. PCs [Ano94-37]. PeerLogic [Ano95-28]. [Ano94-101, Ano95-62]. DB [Ano95-66]. O [Ano95y]. performance [Ano95-29]. Pentium [Ano94-233, Ano94-184]. Pentiums [Ano95n]. people [Ano94-40]. PerfectOffice [Ano94-234]. performance [Ano94-182]. performing [Ano94-46]. personal [Ano95-27]. Pharmaceutical [Ano94-170]. picture [Ano94n, Ano95b]. piece [Ano94n]. Pipes [Ano95-28]. pits [Ano95n]. pizza [Ano94-165]. PL [Ano94-113]. PL/ [Ano94-113]. place [Ano95m]. plan [Ano94-176]. plane [Ano94-114]. planning [Ano94u]. plans [Ano94-70, Ano95-63]. Plastics [Ano94-62]. platform [Ano94-228, Ano94-152]. platforms [Ano94-225, Ano94-111]. Platinum [Ano94-171]. play [Ano95-44]. playtime [Ano95-67]. please [Ano94-228]. plug [Ano95-44]. plug-and-play [Ano95-44]. plus [Ano94a]. point [Ano94-102]. point-and-click [Ano94-102]. Pooh [Ano94-87]. popular [Ano94s]. port [Ano94-113, Ano95i]. portability [Ano94-229]. portable [Ano95-61]. Porting [Ano95-56]. position [Ano94-108]. positioned [Ano94-213]. positive [Ano94e]. possible [Ano94-232]. potential [Ano94-39]. Power [Ano94-168]. PowerBuilder [Ano94-187]. powerful [Ano95e]. PowerPC [Ano94-84, Ano94-63]. PowerPC-based [Ano94-84, Ano94-172]. PPP [Ano94-219]. presentation [Ano94t]. preserve [Ano94-147]. Pretty [Ano94-176]. previous [Ano95n]. price [Ano94-66]. price/performance [Ano95-29]. pricing [Ano95-47]. primacy [Ano94-156]. printer [Ano94-67]. printers [Ano94-145]. Privacy [Ano94-176]. problems [Ano94-136]. process [Ano94-114]. processing [Ano94c, Ano95d]. processor [Ano94-48]. processor-independent [Ano94-48]. processors [Ano94-144]. product [Ano94-168, Ano94-183]. productivity [Ano94-224]. products [Ano94-179, Ano94-41]. professionals [Ano94-40]. program [Ano94-221]. programming [Ano94-86, Ano94-187]. programs [Ano95h]. progress [Ano94-82, Fei94a]. Project [Ano94-47]. proliferating [Ano94-141]. promised [Ano94-168]. promote [Ano94-179]. prompted [Ano94-183]. proof [Ano94-218]. proper [Ano94o]. prosecution [Ano94d]. protocol [Ano94l]. provide [Ano94-136]. providers [Ano94-179]. provides [Ano94q]. providing [Ano94-101]. purchase [Ano94-49]. purpose [Ano95-56]. push [Ano95b]. put [Ano94-224].

QMS [Ano94-145]. quality [Ano94k]. quarter [Ano95-64]. queries [Ano94-203]. query [Ano94-30]. questions [Ano94-166]. quick [Ano94-195]. quite [Ano94-106].

rages [Ano94-232]. RAID [Ano95g]. raising [Ano94-212]. ran [Ano94-168]. range [Ano95-29]. Rap [Lev94a]. rather [Ano95-46]. RDI [Ano95e]. reached [Ano94-150]. reaches [Ano94-193]. read [Ano94-180]. readies [Ano94t]. real [Ano94-109]. receivable [Ano94-107]. receive [Ano94-116]. receiving [Ano94-210]. recognition [Ano94-166]. recordable [Ano94-223]. redefining [Ano94-46]. redesigned [Ano94a]. redundancy [Ano94-69]. redundant [Ano94-76]. refrigerator [Ano94-218]. regulatory [Ano94x]. relate [Ano94o]. relational [Ano94-147]. Release [Ano94f]. released [Ano94-35]. reliability [Ano94-189]. remains [Ano94-141]. remote [Ano94-164]. remote-access [Ano94-164]. remote-connectivity [Ano94-204]. replaced [Ano94-191]. replacement [Ano94-134]. reporting [Ano94q]. represents [Ano94u]. Request [Ano94-138]. requires [Ano94u]. resilient [Ano94-69, Ano95x]. resolution [Ano94-189]. resource [Ano95w]. respectively [Ano94-160]. results [Ano95-64, Ano94-53]. rethinking [Ano94-48]. retool [Ano94-117]. reusable [Ano94u]. reversal [Ano94-177]. review [Ano94-174, Ano94-229]. revs [Ano94-182]. right [Ano94-186]. Ring [Ano95a]. RISC [Ano94-184]. robust [Ano94v]. rocketing [Ano94-131]. role [Ano94p, Ano94-177]. roll [Ano94-184]. rolling [Ano94-225]. rolls [Ano94a]. round [Ano94-69]. round-the-clock [Ano94-69]. rounds [Ano95-64]. route [Ano94-151]. router [Ano94-75, Ano94-75]. routers [Ano94-191, Ano94-211]. routing [Ano94-34]. RPC [Ano94-232]. RPC-vs.-messaging [Ano94-232]. RS [Ano94-84]. RS/ [Ano94-84]. ruled [Ano94x]. run [Ano94-74]. running [Ano94-77, Ano95x]. runs [Ano94j]. rushing [Ano94-211].

sales [Ano94-183]. same [Ano94h]. San [Ano94-31]. Santa [Ano94-215]. SAS [Ano94-178]. says [Ano94-32]. scheduled [Ano94-64]. scheduling [Ano94-28]. Science [Ano95-68]. scripting [Ano94-167]. SCSI [Ano94-132]. second [Ano95x]. Secure [Ano95-58]. secures [Ano94-47]. securing [Ano95-53]. security [Ano94-146]. seek [Ano94-43]. seen [Ano94z]. sees [Ano94-151]. self [Ano94-170]. self-managing [Ano94-170]. sell [Ano94-53]. send [Ano94-180]. senior [Ano95-54]. series [Ano95x]. server [Ano94c]. server-based [Ano95-55]. servers [Ano94-63, Ano94-180]. service [Ano94-179]. services [Ano94-57, Ano94-179]. set [Ano94-112]. sets [Ano95-49]. seven [Ano95-38, Ano94-179]. several [Ano94-178, Ano94-180]. share [Ano95g]. sharing [Ano94-158]. shell [Ano94m]. Shield [Ano95q]. shift [Ano95b]. shine [Ano94-223]. ship [Ano94-188]. shipped [Ano94-160]. shipping [Ano94-78]. shortage [Ano94-39]. should [Ano94-232]. show [Ano95-64, Ano94p]. shows [Ano94-102]. Shrink [Lev94a]. side [Ano95x]. sidelined [Ano94-39]. sight [Ano94l]. Silicon [Ano94-182]. simple [Ano94-186]. simultaneous [Ano94-64]. single [Ano94-228]. single- [Ano94-149]. single-platform [Ano94-228]. sit [Ano95-64]. sites [Ano95l]. slammed [Ano95-68]. SLIP [Ano94m]. slots [Ano94-206]. slow [Ano94-39, Ano94-183]. small [Ano94i]. smaller [Ano94-113]. Smalltalk [Ano94-210]. Smalltalks [Ano95v]. SMS [Ano95w]. SNA [Ano94-74]. SNMP [Ano94-112, Ano94-105]. snoop [Ano94-176]. SoftBench [Ano94-65]. software [Ano94-74, Ano94-115]. Solaris [Ano94-161]. SolarNet [Ano94-198]. solution [Ano95s]. solutions [Ano95-44]. solving [Ano94-136]. some [Ano94-46]. soon [Ano94-207]. sort [Ano94-87]. sounds [Ano94-186]. sources [Ano94-150]. Space [Ano95-68]. SPARC [Ano94-61]. SPARC-based [Ano94-61]. special [Ano94o]. specifically [Ano94-145]. specification [Ano94-172, Ano94-138]. speed [Ano94-34, Ano94-184]. spinoff [Ano94d]. spoke [Ano94-153]. spotlight [Ano94-155]. spreadsheet [Ano94-183]. spring [Ano94-172]. Spry [Ano94-185]. Spyglass [Ano94n]. SQL [Ano94-101, Ano94-150]. SQL*Net [Ano95-53]. SQLWindows [Ano94-187]. staff [Ano94-108]. stage [Ano95-47]. stagestruck [Ano94p]. staging [Ano94-171]. standalone [Ano94-216]. standard [Ano94-101, Ano95m]. standards [Ano94-83, Fei94a]. Stantards [Fei94m]. start [Ano94-39, Ano94-188]. Start-up [Ano94-188]. stealing [Ano94-110]. Stealth [Ano94-114]. Stearns [Ano94-68]. sticks [Ano94-230]. stiffens [Ano94-177]. STL [Ano95-61]. Stop [Ano94-27, Ano94-191]. storage [Ano94-76]. store [Ano94-231]. strange [Ano94-87]. strategic [Ano94-208]. strategies [Ano94-137]. strategy [Ano95w]. strengthen [Ano94-103]. styles [Ano95-31]. success [Ano94-41]. suite [Ano94-234]. suited [Ano94-223]. summer [Ano94-203]. Sun [Ano94s, Ano94-195]. SunPC [Ano95j]. SunSoft [Ano95j, Ano94-197]. SuperStack [Ano95a]. support [Ano94-101, Ano94-210]. supporting [Ano94p]. supports [Ano94-178]. sure [Ano94-228, Ano94-201]. survey [Ano94-53]. Swiss [Ano94-202]. switch [Ano95i, Ano94-119]. switches [Ano95r, Ano94-191]. switching [Ano94-211]. Sybase [Ano94-77, Ano94-203]. Symantec [Ano94-204]. symmetric [Ano94-81]. symmetrical [Ano94-84]. system [Ano94-222, Ano95-28]. Systems [Hun94a].

T [Ano94-104]. T/ [Ano94-37]. tackles [Ano95w]. take [Ano94-155]. takes [Ano94-229]. taking [Ano94-223]. talk [Ano94-230]. talks [Ano94e]. TCP [Ano94h, Ano95-62]. TCP/ [Ano94h, Ano95-62]. team [Ano94-104]. teams [Ano94-28]. tech [Ano94e]. technical [Ano94-191]. technologies [Ano94o, Ano94-32]. technology [Ano94-211, Ano94-39]. Teknekron [Ano94-140]. Tektronix [Ano94-206]. Telecom [Ano94-151]. telecommunications [Ano94-104]. telephones [Ano94-59]. telephony [Ano94-37]. Telescope [Ano95-68, Ano95-68]. terminals [Ano94-206, Ano94-71]. testing [Ano94-161]. tests [Ano94-207]. Texaco [Ano94-152]. Texas [Ano95w]. thanks [Ano94j]. their [Ano94-168]. them [Ano95n, Ano94-141]. there [Ano94-141, Ano94-106]. these [Ano94-137]. they’re [Ano94-53]. things [Ano94d]. third [Ano94-234, Ano95-63]. third-party [Ano94-234, Ano95-63]. though [Ano94-172]. threatens [Ano94-211]. three [Ano94-208]. throughput [Ano94-115]. thunder [Ano94-110]. ticketing [Ano95z]. time [Ano95j]. tips [Ano94r]. today [Ano94-157]. together [Ano94l]. Token [Ano95a]. tons [Ano95-68]. too [Ano94z]. tool [Ano94-65, Ano94-102]. toolkit [Ano94-228]. tools [Ano94-222, Ano94t]. toolset [Ano94q]. toolsets [Ano95-49]. Toolst [Ano94-229]. top [Ano94-46]. top-performing [Ano94-46]. tour [Ano94-191]. TPC [Ano95-64]. track [Ano94-186]. trade [Ano94-64]. traffic [Ano94-34]. transaction [Ano94c]. transaction-processing [Ano94c]. transforms [Ano94-205]. translation [Ano94-196]. translations [Ano94-87]. transmissions [Ano95-58]. transports [Ano94-180]. tricks [Ano94-132]. Trinzic [Ano95-65]. trivial [Ano94x]. trouble [Ano95z]. trying [Ano94-131]. tuned [Ano95y]. turned [Ano95q]. turnkey [Ano94-195]. turns [Ano94m]. Tuxedo [Ano94-227]. TV [Ano94k]. TV-quality [Ano94k]. twice [Ano94-184]. two [Ano94-226, Ano94-219]. two-part [Ano95x]. types [Ano94-166].

understand [Ano94-232]. understanding [Ano94-137]. Uniface [Ano94-187]. Unify [Ano94-220]. unique [Ano95-32]. Universal [Ano94-208].

Unix [Ano94-197]. Unix-based [Ano95o]. Unix-on-Intel [Ano95-47]. UnixWare [Ano95-47, Ano94-221]. until [Ano94-70]. unveil [Ano94-195]. unveiled [Ano94-76]. unveils [Ano94y]. upcoming [Ano94-118]. updates [Ano94-171]. upgrade [Ano94s]. upgraded [Ano95-50]. upgrades [Ano94-197]. uptime [Ano94-69]. US$35 [Ano95-29]. US$50 [Ano95-29]. US$7 [Ano94-67]. use [Ano94-221, Ano94-146]. used [Ano94-36]. user [Ano94r, Atk94g]. users [Ano94-234, Ano94-222]. uses [Ano94-57]. using [Ano94-63]. utilities [Ano94-197]. UX [Ano95y].

V.3 [Ano94-174]. vague [Ano94-54]. value [Ano94-48]. ValuePoint [Ano94-83]. Vantive [Ano94-35]. variety [Ano95-65]. vast [Ano94-114]. vastly [Ano94-115]. VBXs [Ano94-134]. vendor [Ano94-180]. vendors [Ano94-166, Ano94-225]. version [Ano94-233, Ano94-226]. versions [Ano94-160]. vertical [Ano94-235]. via [Ano94-112]. video [Ano94-104]. video-conferencing [Ano94-104]. videoconferencing [Ano94k]. view [Ano94-114, Azz94e]. views [Ano94q, Ano94-142]. Village [Ano95c]. Vines [Ano95p]. virtual [Ano95r]. Visix [Ano94-228]. visual [Ano94-220, Ano94-229]. VisualWorks [Ano95-43]. vital [Ano94-50, Ano94-201]. voice [Ano94-166]. voiced [Ano94-105]. voltage [Ano94g]. Voyager [Ano95e]. vs [Ano94-232].

Wabi [Ano94-199]. wacky [Ano95-38, Ano95-67]. waiting [Ano94-233]. wants [Ano94-106]. warehouses [Ano94-203]. warehousing [Ano94-41, Ano94-101]. wares [Ano94-164]. warm [Ano94-58]. Warp [Ano94-78]. was [Ano94-168]. Washington [Ano94-205]. Watch [Ano94-88]. watching [Ano95-68]. way [Ano94-106, Ano94-230]. ways [Ano94-192]. weary [Ano95-67]. weather [Ano94-230]. Web [Ano94-36, Ano95-52]. week [Ano94-164]. well [Ano94-223]. were [Ano94x]. western [Ano94p]. where [Ano94-192]. whether [Ano94-191]. which [Ano94y]. while [Ano94-46, Ano94-232]. Whitetree [Ano95i]. who [Ano94-40]. wide [Ano94-192]. wide-area [Ano94-192]. widely [Ano94-36]. widgets [Ano95-30]. wild [Ano95-38]. will [Ano94-234]. win [Ano94-152]. Windows [Ano94-168, Ano94z]. Windows-based [Ano94-168]. Windows-to-Unix [Ano94-226]. winners [Ano94-49]. Winsock [Ano95-50]. wireless [Ano94-39]. withdrawn [Ano94x]. within [Ano94-216]. wolf [Ano94-218]. wolf-proof [Ano94-218]. word [Ano94-184]. word- [Ano95-31]. WordPerfect [Ano94-160, Ano94-234]. words [Ano94-156]. work [Ano94-132]. workflow [Ano94-51, Ano94-235]. workgroup [Ano94-152]. workgroup-computing [Ano94-152]. working [Ano94r]. Workshop [Ano95-32]. workstation [Ano94-182]. workstations [Ano94-53]. World [Ano94-36]. would [Ano94-166]. writing [Ano94-201].

X [Ano94-206]. X.400 [Ano94-55]. X.500 [Ano94-119]. XP [Ano95-50].

year  [Ano95-68,  Ano94-70].

years [Ano95-54]. You’ll [Ano94-54].

References

Anonymous:1994:RRH

[Ano94a] Anonymous. 35om rolls out a redesigned hub chassis, plus a multiprocessor Ethernet adapter. Open Systems Today, 155:32–??, August 1994. ISSN 1061-0839.

Anonymous:1994:ATA

[Ano94b] Anonymous. 3Com adds two ASIC-based midrange switching hub modules for Ethernet environments. Open Systems Today, 165:12–??, December 1994. ISSN 1061-0839.

Anonymous:1994:DAA

[Ano94c] Anonymous. A DEC Alpha AXP server has edged out an IBM mainframe in the latest transaction-processing benchmark competition. Open Systems Today, 153:46–??, July 1994. ISSN 1061-0839.

Anonymous:1994:HSP

[Ano94d] Anonymous. A Honeywell spinoff plans to unveil an Internet firewall server that, among other things, can help identify intruders for prosecution. Open Systems Today, 161:4–??, October 1994. ISSN 1061-0839.

Anonymous:1994:LFR

[Ano94e] Anonymous. A law firm that represents high-tech companies talks about its positive Internet experience. Open Systems Today, 164:64–??, November 1994. ISSN 1061-0839.

Anonymous:1994:LAR

[Ano94f] Anonymous. A look at Release 6 of JYACC’s JAM multiplatform development package. Open Systems Today, 154:61–??, July 1994. ISSN 1061-0839.

Anonymous:1994:LPW

[Ano94g] Anonymous. A low-voltage 75-MHz Pentium will run cool enough to be used in notebook computers. Open Systems Today, 161:6–??, October 1994. ISSN 1061-0839.

Anonymous:1994:NIM

[Ano94h] Anonymous. A new Internet mail package gives mobile users the same TCP/ IP interface as LAN users. Open Systems Today, 154: 32–??, July 1994. ISSN 1061-0839.

Anonymous:1994:NLR

[Ano94i] Anonymous. A new low-end router is designed to connect small companies’ Ethernet nets to the Internet. Open Systems Today, 160:46–??, October 1994. ISSN 1061-0839.

Anonymous:1994:NPM

[Ano94j] Anonymous. A new Power Macintosh runs DOS and Windows apps, thanks to a 486 coprocessor card. Open Systems Today, 164:8–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NTV

[Ano94k] Anonymous. A new TV-quality videoconferencing system from Avistar runs on its own network and doesn’t interfere with the corporate LAN. Open Systems Today, 165:30–??, December 1994. ISSN 1061-0839.

Anonymous:1994:NIP

[Ano94l] Anonymous. A next-generation Internet protocol comes into sight at an engineers’-get-together. Open Systems Today, 155:4–??, August 1994. ISSN 1061-0839.

Anonymous:1994:PPM

[Ano94m] Anonymous. A package for PCs and Macs turns Internet shell accounts into SLIP, giving users the full benefits of Net connection. Open Systems Today, 162: 58–??, October 1994. ISSN 1061-0839.

Anonymous:1994:PMP

[Ano94n] Anonymous. A piece of the Mosaic picture comes into focus as Spyglass obtains a master license to the code from NCSA. Open Systems Today, 157:8–??, September 1994. ISSN 1061-0839.

Anonymous:1994:PCO

[Ano94o] Anonymous. A proper comparison of OpenDoc and OLE means considering not just the base technologies, but how these technologies relate to a user’s special needs. Open Systems Today, 155: 28–??, August 1994. ISSN 1061-0839.

Anonymous:1994:SKL

[Ano94p] Anonymous. A stagestruck Karl lands a supporting role in a oneman country-and-western show. Open Systems Today, 158:90–??, September 1994. ISSN 1061-0839.

Anonymous:1994:TTE

[Ano94q] Anonymous. A third-party toolset that enhances OpenView provides reporting capabilities, custom network views, and more. Open Systems Today, 157:52–??, September 1994. ISSN 1061-0839.

Anonymous:1994:UOT

[Ano94r] Anonymous. A user offers tips on working with the Distributed Computing Environment. Open Systems Today, 153:52–??, July 1994. ISSN 1061-0839.

Anonymous:1994:APR

[Ano94s] Anonymous. Adobe plans to release Illustrator 5.5 for Sun in November, bringing users a major upgrade for the popular graphics package. Open Systems Today, 158:34–??, September 1994. ISSN 1061-0839.

Anonymous:1994:ARA

[Ano94t] Anonymous. Adobe readies Acrobat 2.0, a long-awaited version of its document presentation software with enhanced integration tools. Open Systems Today, 158:5–??, September 1994. ISSN 1061-0839.

Anonymous:1994:AIR

[Ano94u] Anonymous. Although it represents much of the allure of objects, creating reusable code also requires some planning. Open Systems Today, 155:60–??, August 1994. ISSN 1061-0839.

Anonymous:1994:EVC

[Ano94v] Anonymous. An eight-member vendor coalition is offering a new, robust messaging infrastructure. Open Systems Today, 165:4–??, December 1994. ISSN 1061-0839.

Anonymous:1994:ILA

[Ano94w] Anonymous. An inside look at how Burlington Coat Factory plans to market its products on the Internet. Open Systems Today, 156:66–??, August 1994. ISSN 1061-0839.

Anonymous:1994:OAW

[Ano94x] Anonymous. An Oracle advertisement was ruled misleading by a British regulatory body; the ad was withdrawn, but Oracle maintains that the offenses were trivial. Open Systems Today, 162: 8–??, October 1994. ISSN 1061-0839.

Anonymous:1994:AUM

[Ano94y] Anonymous. Apple unveils the Mac System 7.5 OS, which will come with new PC compatibility and networking features. Open Systems Today, 154:6–??, July 1994. ISSN 1061-0839.

Anonymous:1994:ADL

[Ano94z] Anonymous. Apple’s decision to license its System 7 operating system is seen as a defensive move that may be too little, too late to challenge Windows. Open Systems Today, 159:5–??, September 1994. ISSN 1061-0839.

Anonymous:1994:ASO

[Ano94-27] Anonymous. At Stop 8, OST’s 11-part technical tour looks at on-line transaction processing, and how it is moving closer to end users. Open Systems Today, 158:48–??, September 1994. ISSN 1061-0839.

Anonymous:1994:BTC

[Ano94-28] Anonymous. Banyan teams with Collabra to add conferencing and scheduling to its mail platforms. Open Systems Today, 161:24–??, October 1994. ISSN 1061-0839.

Anonymous:1994:BUA

[Ano94-29] Anonymous. Beta users applaud NetWare 4.1, citing enhanced NDS and new TCP/ IP implementation. Open Systems Today, 163:5–??, November 1994. ISSN 1061-0839.

Anonymous:1994:BOU

[Ano94-30] Anonymous. Business Objects unveiled an upgrade of its query tool with improved reporting, multilevel alerts, and off-line operation. Open Systems Today, 165:8–??, December 1994. ISSN 1061-0839.

Anonymous:1994:CDP

[Ano94-31] Anonymous. Cellular Digital Packet Data service comes to the San Francisco Bay area, but the once-hyped technology has been slow to catch on. Open Systems Today, 162:36–??, October 1994. ISSN 1061-0839.

Anonymous:1994:CBF

[Ano94-32] Anonymous. CEO Bob Frankenberg says Novell will focus on its core technologies. Open Systems Today, 161:30–??, October 1994. ISSN 1061-0839.

Anonymous:1994:CMB

[Ano94-33] Anonymous. Chase Manhattan Bank looked at ATM, but has backed off, at least for now. Open Systems Today, 154:1–??, July 1994. ISSN 1061-0839.

Anonymous:1994:CII

[Ano94-34] Anonymous. Cisco is introducing routing software products that speed SNA traffic on TCP/ IP networks. Open Systems Today, 163:4–??, November 1994. ISSN 1061-0839.

Anonymous:1994:CVE

[Ano94-35] Anonymous. Clarify and Vantive have each released multiplatform versions of their customer service applications. Open Systems Today, 153:24–??, July 1994. ISSN 1061-0839.

Anonymous:1994:CAN

[Ano94-36] Anonymous. Commercial alternatives to NCSA’s Mosaic World Wide Web browser compete with the widely used freeware. Open Systems Today, 158:68–??, September 1994. ISSN 1061-0839.

Anonymous:1994:CAG

[Ano94-37] Anonymous. Compaq, AT and T/ GIS, and IBM join a growing number of vendors that are offering PCs with built-in telephony capabilities. Open Systems Today, 159:42–??, September 1994. ISSN 1061-0839.

Anonymous:1994:CDG

[Ano94-38] Anonymous. Connectivity device gives PC and Mac users low-cost, 500-Kbps Internet link. Open Systems Today, 163:51–??, November 1994. ISSN 1061-0839.

Anonymous:1994:CIS

[Ano94-39] Anonymous. COPD is off to a slow start, as a shortage of apps has sidelined many potential users of the wireless datacom technology. Open Systems Today, 158:40–??, September 1994. ISSN 1061-0839.

Anonymous:1994:DAD

[Ano94-40] Anonymous. Data access from the desktop is more and more the job of people who don’t know much about computers, so the job of IS professionals is to make that access easy. Open Systems Today, 159:48–??, September 1994. ISSN 1061-0839.

Anonymous:1994:DAI

[Ano94-41] Anonymous. Data access is the key to success as OST looks at data warehousing products. Open Systems Today, 160:58–??, October 1994. ISSN 1061-0839.

Anonymous:1994:DGE

[Ano94-42] Anonymous. Data General expands its AViiON line of storage products with a CD jukebox, management software, and more. Open Systems Today, 159: 44–??, September 1994. ISSN 1061-0839.

Anonymous:1994:DWIa

[Ano94-43] Anonymous. Data warehousing isn’t just for archiving anymore, as users seek to gain a competitive edge from the technology. Open Systems Today, 160:82–??, October 1994. ISSN 1061-0839.

Anonymous:1994:DNE

[Ano94-44] Anonymous. Dawn of a new era?

Open Systems Today, 155:1–??, August 1994. ISSN 1061-0839.

Anonymous:1994:DRN

[Ano94-45] Anonymous. DCA rolls out a new enterprise software tool that manages software distribution, application metering, and remote diagnosis. Open Systems Today, 158:56–??, September 1994. ISSN 1061-0839.

Anonymous:1994:DDS

[Ano94-46] Anonymous. DEC debuts some top-performing machines while redefining its organizational focus. Open Systems Today, 155: 44–??, August 1994. ISSN 1061-0839.

Anonymous:1994:DDE

[Ano94-47] Anonymous. DECathena, Digital Equipment’s version of MIT’s Project Athena, managers and secures huge networks with aplomb. Open Systems Today, 160:76–??, October 1994. ISSN 1061-0839.

Anonymous:1994:DEA

[Ano94-48] Anonymous. Digital Equipment apparently is rethinking the value of processor-independent NetWare. Open Systems Today, 157:5–??, September 1994. ISSN 1061-0839.

Anonymous:1994:DEU

[Ano94-49] Anonymous. Digital Equipment users look like the big winners in Oracle’s purchase of DEC’s DMBS lines. Open Systems Today, 159:35–??, September 1994. ISSN 1061-0839.

Anonymous:1994:DJS

[Ano94-50] Anonymous. Distributed job scheduling tools are vital to today’s heterogeneous computing. Open Systems Today, 165:56–??, December 1994. ISSN 1061-0839.

Anonymous:1994:DWIb

[Ano94-51] Anonymous. Documentum 2.0 will include workflow features and a better management-system interface. Open Systems Today, 164:20–??, November 1994. ISSN 1061-0839.

Anonymous:1994:EUB

[Ano94-52] Anonymous. Early users of Borland’s dBase for Windows give it high marks, despite its lack of interoperability with other database environments. Open Systems Today, 159:4–??, September 1994. ISSN 1061-0839.

Anonymous:1994:EWS

[Ano94-53] Anonymous. Entry-level workstations sell better when they’re loaded with options, according to recent survey results. Open Systems Today, 158:62–??, September 1994. ISSN 1061-0839.

Anonymous:1994:EFA

[Ano94-54] Anonymous. Ever fret about late software or vague manuals? You’ll love Karl’s new friends. Open Systems Today, 154:96–??, July 1994. ISSN 1061-0839.

Anonymous:1994:FLE

[Ano94-55] Anonymous. Faced with a lack of E-mail connectivity, the Energy Department chose an X.400 gateway. Open Systems Today, 156: 1–??, August 1994. ISSN 1061-0839.

Anonymous:1994:FAA

[Ano94-56] Anonymous. Financial and accounting applications add ODBC support as a way to access SQL databases. Open Systems Today, 162:28–??, October 1994. ISSN 1061-0839.

Anonymous:1994:FSG

[Ano94-57] Anonymous. Financial services giant Fidelity Investments uses distributed object architecture to build an enterprise-wide open environment. Open Systems Today, 155:62–??, August 1994. ISSN 1061-0839.

Anonymous:1994:FNN

[Ano94-58] Anonymous. Firefox’s NOV’IX NetWare gateway Version 2.2 will be released this week, adding FDDI support and warm failover capability. Open Systems Today, 165:47–??, December 1994. ISSN 1061-0839.

Anonymous:1994:FIL

[Ano94-59] Anonymous. Four industry leaders have formed a group to promote interoperability standards for integrating computers and telephones. Open Systems Today, 165:38–??, December 1994. ISSN 1061-0839.

Anonymous:1994:FVW

[Ano94-60] Anonymous. Four vendors will link their CORBA implementations to Windows OLE 2.0. Open Systems Today, 156:43–??, August 1994. ISSN 1061-0839.

Anonymous:1994:FSN

[Ano94-61] Anonymous. Fujitsu’s SPARCbased NFS server integrates several storage devices in one box. Open Systems Today, 161:39–??, October 1994. ISSN 1061-0839.

Anonymous:1994:GPC

[Ano94-62] Anonymous. GE Plastics customers find a home on the Net: a World Wide Web address where they can access 1,500 pages of technical data. Open Systems Today, 165:72–??, December 1994. ISSN 1061-0839.

Anonymous:1994:GBI

[Ano94-63] Anonymous. Groupe Bull has introduced its Escala servers, which are the industry’s first symmetric multiprocessor systems using the PowerPC. Open Systems Today, 159:6–??, September 1994. ISSN 1061-0839.

Anonymous:1994:HCL

[Ano94-64] Anonymous. Here come lots of new DBMS technology, to be exhibited at two simultaneous trade shows scheduled for next month. Open Systems Today, 164:5–??, November 1994. ISSN 1061-0839.

Anonymous:1994:HID

[Ano94-65] Anonymous. Hewlett–Packard introduces distributed versions of SoftBench development-tool framework. Open Systems Today, 158:54–??, September 1994. ISSN 1061-0839.

Anonymous:1994:HOH

[Ano94-66] Anonymous. Hewlett–Packard offers HP 9000 price cuts to entice users to migrate from AS/ 400s; IBM counters with marketing, price matching. Open Systems Today, 160:5–??, October 1994. ISSN 1061-0839.

Anonymous:1994:HUC

[Ano94-67] Anonymous. Hewlett–Packard unveils a color laser printer at a market-leading price of US$7,295. Open Systems Today, 160:87–??, October 1994. ISSN 1061-0839.

Anonymous:1994:HIB

[Ano94-68] Anonymous. How investment banker Bear Stearns got a handle on power faxing. Open Systems Today, 154:71–??, July 1994. ISSN 1061-0839.

Anonymous:1994:HDF

[Ano94-69] Anonymous. How to design fault-resilient systems, which offer nearly the same round-the-clock system uptime as hardware redundancy — for a lot less money. Open Systems Today, 165:62–??, December 1994. ISSN 1061-0839.

Anonymous:1994:HPD

[Ano94-70] Anonymous. HP plans distributed capabilities for Open-View, but not until the middle of next year. Open Systems Today, 156:4–??, August 1994. ISSN 1061-0839.

Anonymous:1994:HDS

[Ano94-71] Anonymous. Human Designed Systems is one of the first companies to incorporate audio and video in a line of X terminals. Open Systems Today, 161:42–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IMAb

[Ano94-72] Anonymous. IBM and Mas-Par announced high-performance data warehousing packages. Open Systems Today, 165:6–??, December 1994. ISSN 1061-0839.

Anonymous:1994:IED

[Ano94-73] Anonymous. IBM expands database support in an AIX NetView release to be announced next month. Open Systems Today, 156:6–??, August 1994. ISSN 1061-0839.

Anonymous:1994:IEA

[Ano94-74] Anonymous. IBM has enhanced its AnyNet products, adding software that allows NetBIOS apps to run over SNA networks, and more. Open Systems Today, 153:26–??, July 1994. ISSN 1061-0839.

Anonymous:1994:IIR

[Ano94-75] Anonymous. IBM has improved its 6611 router models, and will be adding a new low-end router. Open Systems Today, 159:30–??, September 1994. ISSN 1061-0839.

Anonymous:1994:IUN

[Ano94-76] Anonymous. IBM has unveiled its new generation of redundant storage devices for mainframes. Open Systems Today, 153:48–??, July 1994. ISSN 1061-0839.

Anonymous:1994:IISa

[Ano94-77] Anonymous. IBM is set to announce a tool for managing performance of Oracle, Sybase, and DB2 databases running on the same network. Open Systems Today, 160:4–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IISb

[Ano94-78] Anonymous. IBM is shipping OS/ 2 3.0, dubbed OS/ 2 Warp, billing it as a “here today” alternative to Microsoft’s forthcoming Windows 95. Open Systems Today, 162:52–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IPD

[Ano94-79] Anonymous. IBM plans to deliver a distributed version of Smalltalk next year. Open Systems Today, 163:10–??, November 1994. ISSN 1061-0839.

Anonymous:1994:IWU

[Ano94-80] Anonymous. IBM will use its own APIs in the forthcoming “Lakes” collaborative environment. Open Systems Today, 161: 4–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IFS

[Ano94-81] Anonymous. IBM’s first symmetric multiprocessing systems for Unix are due out this month. Open Systems Today, 157:4–??, September 1994. ISSN 1061-0839.

Anonymous:1994:INO

[Ano94-82] Anonymous. IBM’s new OS/ 400 release shows great progress toward openness. Open Systems Today, 154:62–??, July 1994. ISSN 1061-0839.

Anonymous:1994:INPb

[Ano94-83] Anonymous. IBM’s new Pentium PCs and servers adhere more broadly to industry standards than its existing PC and ValuePoint models. Open Systems Today, 162:54–??, October 1994. ISSN 1061-0839.

Anonymous:1994:INPa

[Ano94-84] Anonymous. IBM’s new PowerPC-based RS/ 6000s include a low-end workstation with PC features, and symmetrical multiprocessing servers. Open Systems Today, 161:5–??, October 1994. ISSN 1061-0839.

Anonymous:1994:ISC

[Ano94-85] Anonymous. IBM’s support for C++ and object technology on MVS mainframes leads Unix Expo news. Open Systems Today, 162:46–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IPT

[Ano94-86] Anonymous. Icon-based programming tools make it easy for both corporate developers and end users to customize PC databases. Open Systems Today, 160: 36–??, October 1994. ISSN 1061-0839.

Anonymous:1994:WPE

[Ano94-87] Anonymous. In which Pooh encounters a computer’s strange translations into French and, sort of, back. Open Systems Today, 161:66–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IWa

[Ano94-88] Anonymous. Industry Watch.

Open Systems Today, 153:17–??, July 1994. ISSN 1061-0839.

Anonymous:1994:IWb

[Ano94-89] Anonymous. Industry Watch.

Open Systems Today, 154:25–??, July 1994. ISSN 1061-0839.

Anonymous:1994:IWc

[Ano94-90] Anonymous. Industry Watch.

Open Systems Today, 155:25–??, August 1994. ISSN 1061-0839.

Anonymous:1994:IWd

[Ano94-91] Anonymous. Industry Watch.

Open Systems Today, 156:22–??, August 1994. ISSN 1061-0839.

Anonymous:1994:IWe

[Ano94-92] Anonymous. Industry Watch.

Open Systems Today, 157:24–??, September 1994. ISSN 1061-0839.

Anonymous:1994:IWf

[Ano94-93] Anonymous. Industry Watch.

Open Systems Today, 158:30–??, September 1994. ISSN 1061-0839.

Anonymous:1994:IWg

[Ano94-94] Anonymous. Industry Watch. Open Systems Today, 159:16–??, September 1994. ISSN 1061-0839.

Anonymous:1994:IWh

[Ano94-95] Anonymous. Industry Watch. Open Systems Today, 160:28–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IWi

[Ano94-96] Anonymous. Industry Watch. Open Systems Today, 161:19–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IWj

[Ano94-97] Anonymous. Industry Watch. Open Systems Today, 162:24–??, October 1994. ISSN 1061-0839.

Anonymous:1994:IWk

[Ano94-98] Anonymous. Industry Watch. Open Systems Today, 163:24–??, November 1994. ISSN 1061-0839.

Anonymous:1994:IWl

[Ano94-99] Anonymous. Industry Watch. Open Systems Today, 164:14–??, November 1994. ISSN 1061-0839.

Anonymous:1994:IWm

[Ano94-100] Anonymous. Industry Watch. Open Systems Today, 165:24–??, December 1994. ISSN 1061-0839.

Anonymous:1994:IBE

[Ano94-101] Anonymous. Information Builders’ EDA/ SQL middleware product line has been enhanced and overhauled, providing more connectivity options, support for standard client APIs and data manipulation features for warehousing. Open Systems Today, 153:8–??, July 1994. ISSN 1061-0839.

Anonymous:1994:ISN

[Ano94-102] Anonymous. Informix shows NewEra, its object-oriented point-and-click development tool. Open Systems Today, 154:5–??, July 1994. ISSN 1061-0839.

Anonymous:1994:IWS

[Ano94-103] Anonymous. Informix will strengthen its platform independence this fall with DCE connectivity and more. Open Systems Today, 156:5–??, August 1994. ISSN 1061-0839.

Anonymous:1994:IAT

[Ano94-104] Anonymous. Intel and AT and T team up to offer videoconferencing services over AT and T’s telecommunications network. Open Systems Today, 157: 30–??, September 1994. ISSN 1061-0839.

Anonymous:1994:IMAa

[Ano94-105] Anonymous. IS managers at a recent conference voiced concern over having to support two management standards: DMI and SNMP. Open Systems Today, 164:4–??, November 1994. ISSN 1061-0839.

Anonymous:1994:IAM

[Ano94-106] Anonymous. It appears that Microsoft wants to be top dog in enterprise computing, but it has quite a way to go to get there. Open Systems Today, 156:26–??, August 1994. ISSN 1061-0839.

Anonymous:1994:KGC

[Ano94-107] Anonymous. Karl gets a call from a desperate deadbeat — and leverages his accounts receivable. Open Systems Today, 159: 66–??, September 1994. ISSN 1061-0839.

Anonymous:1994:KMC

[Ano94-108] Anonymous. Karl must choose between freedom and Santa’s offer of a cozy staff position. Open Systems Today, 165:90–??, December 1994. ISSN 1061-0839.

Anonymous:1994:KNC

[Ano94-109] Anonymous. Karl’s neighbor crafts a network-in-a-box that’s a real antique — inside and out. Open Systems Today, 156:82–??, August 1994. ISSN 1061-0839.

Anonymous:1994:LSS

[Ano94-110] Anonymous. LAN servers are stealing some of the AS/ 400’s thunder in the midrange. Open Systems Today, 161:38–??, October 1994. ISSN 1061-0839.

Anonymous:1994:LIF

[Ano94-111] Anonymous. Legent has introduced a four-module data mover that passes data efficiently between disparate databases and platforms. Open Systems Today, 153:39–??, July 1994. ISSN 1061-

Anonymous:1994:LISa

[Ano94-112] Anonymous. Legent has introduced a set of agent-based products that let users manage databases and Unix systems via SNMP consoles. Open Systems Today, 159:40–??, September 1994. ISSN 1061-0839.

Anonymous:1994:LISb

[Ano94-113] Anonymous. Liant is shipping a new release of PL/ 1 for Unix and Windows NT that will make it easier to port apps to smaller platforms. Open Systems Today, 165:5–??, December 1994. ISSN 1061-0839.

Anonymous:1994:LIU

[Ano94-114] Anonymous. Lockheed is using a process configuration management system that provides a dynamic view of the vast amount of software that’s being developed for the F-22 Stealth fighter plane. Open Systems Today, 159: 54–??, September 1994. ISSN 1061-0839.

Anonymous:1994:LOV

[Ano94-115] Anonymous. Looking to offer vastly improved throughput, NetManaged adds native ISDN support to its client Internet software. Open Systems Today, 157: 38–??, September 1994. ISSN 1061-0839.

Anonymous:1994:LAP

[Ano94-116] Anonymous. Lotus adds a pager interface to cc:Mail, letting remote users send and receive messages up to 240 characters long.

Open Systems Today, 158:36–??, September 1994. ISSN 1061-0839.

Anonymous:1994:LPR

[Ano94-117] Anonymous. Lotus plans to retool its communications server mail engine and add MAPI support. Open Systems Today, 165: 4–??, December 1994. ISSN 1061-0839.

Anonymous:1994:LUC

[Ano94-118] Anonymous. Lotus’ upcoming cc:Mail View will give management capabilities to cc:Mail. Open Systems Today, 162:32–??, October 1994. ISSN 1061-0839.

Anonymous:1994:LWI

[Ano94-119] Anonymous. Lotus will integrate X.500 directory services into its EMX network backbone switch. Open Systems Today, 163:4–??, November 1994. ISSN 1061-0839.

Anonymous:1994:Ma

[Ano94-120] Anonymous. Mail. Open Systems Today, 153:20–??, July 1994. ISSN 1061-0839.

Anonymous:1994:Mb

[Ano94-121] Anonymous. Mail. Open Systems Today, 154:26–??, July 1994. ISSN 1061-0839.

Anonymous:1994:Mc

[Ano94-122] Anonymous. Mail. Open Systems Today, 155:26–??, August 1994. ISSN 1061-0839.

Anonymous:1994:Md

[Ano94-123] Anonymous. Mail. Open Systems Today, 157:26–??, September 1994. ISSN 1061-0839.

Anonymous:1994:Me

[Ano94-124] Anonymous. Mail. Open Systems Today, 158:32–??, September 1994. ISSN 1061-0839.

Anonymous:1994:Mf

[Ano94-125] Anonymous. Mail. Open Systems Today, 159:20–??, September 1994. ISSN 1061-0839.

Anonymous:1994:Mg

[Ano94-126] Anonymous. Mail. Open Systems Today, 161:21–??, October 1994. ISSN 1061-0839.

Anonymous:1994:Mh

[Ano94-127] Anonymous. Mail. Open Systems Today, 162:26–??, October 1994. ISSN 1061-0839.

Anonymous:1994:Mi

[Ano94-128] Anonymous. Mail. Open Systems Today, 163:28–??, November 1994. ISSN 1061-0839.

Anonymous:1994:Mj

[Ano94-129] Anonymous. Mail. Open Systems Today, 164:16–??, November 1994. ISSN 1061-0839.

Anonymous:1994:Mk

[Ano94-130] Anonymous. Mail. Open Systems Today, 165:28–??, December 1994. ISSN 1061-0839.

Anonymous:1994:MVE

[Ano94-131] Anonymous. Major vendors of Email are jazzing up their products, trying to keep up with a market that’s rocketing toward multimedia. Open Systems Today, 156:34–??, August 1994. ISSN 1061-0839.

Anonymous:1994:MSD

[Ano94-132] Anonymous. Making SCSI devices work in harmony often requires a bag of tricks. Open Systems Today, 157:60–??, September 1994. ISSN 1061-0839.

Anonymous:1994:MPP

[Ano94-133] Anonymous. Massively parallel processing is making inroads in commercial applications. Open Systems Today, 157:53–??, September 1994. ISSN 1061-0839.

Anonymous:1994:MID

[Ano94-134] Anonymous. Microsoft is dropping support for VBXs, and the replacement lets developers build compound applications with OLE. Open Systems Today, 163:72–??, November 1994. ISSN 1061-0839.

Anonymous:1994:MWI

[Ano94-135] Anonymous. Microsoft will include native support for Internet multimedia and mail in its Exchange Server messaging engine. Open Systems Today, 154:4–??, July 1994. ISSN 1061-0839.

Anonymous:1994:MIS

[Ano94-136] Anonymous. Middleware is solving the problems of many organizations that must provide users with access to diverse databases and apps. Open Systems Today, 157:64–??, September 1994. ISSN 1061-0839.

Anonymous:1994:MMD

[Ano94-137] Anonymous. More and more these days divining the object strategies of major computer companies means understanding the companies’ partnerships. Open Systems Today, 155: 1–??, August 1994. ISSN 1061-0839.

Anonymous:1994:MVO

[Ano94-138] Anonymous. More vendors are offering products that meet the Common Object Request Broker Architecture specification. Open Systems Today, 155:54–??, August 1994. ISSN 1061-0839.

Anonymous:1994:MCW

[Ano94-139] Anonymous. Mosaic Communications’ World Wide Web browser is now freeware for many users. Open Systems Today, 162: 43–??, October 1994. ISSN 1061-0839.

Anonymous:1994:MBC

[Ano94-140] Anonymous. Moving beyond the consulting business, Teknekron is introducing a packaged version of its distributed architecture. Open Systems Today, 162:8–??, October 1994. ISSN 1061-0839.

Anonymous:1994:MDP

[Ano94-141] Anonymous. Multiprocessor desktops are proliferating, even though there remains a shortage of multithreaded applications for them. Open Systems Today, 165:60–??, December 1994. ISSN 1061-0839.

Anonymous:1994:NV

[Ano94-142] Anonymous. Net Views. Open Systems Today, 156:24–??, August 1994. ISSN 1061-0839.

Anonymous:1994:N

[Ano94-143] Anonymous. NetViews. Open Systems Today, 160:32–??, October 1994. ISSN 1061-0839.

Anonymous:1994:NMN

[Ano94-144] Anonymous. New machines from NEC, AST and Intergraph feature multiple processors and multitasking for the desktop. Open Systems Today, 154:50–??, July 1994. ISSN 1061-0839.

Anonymous:1994:NPD

[Ano94-145] Anonymous. New printers from DEC and QMS are designed specifically for network duty. Open Systems Today, 154:52–??, July 1994. ISSN 1061-0839.

Anonymous:1994:NPA

[Ano94-146] Anonymous. New products are aimed at increasing Internet security and encouraging commercial use. Open Systems Today, 156:8–??, August 1994. ISSN 1061-0839.

Anonymous:1994:NPL

[Ano94-147] Anonymous. New products linking object and relational data preserve investments in legacy databases while adding advantages of object technology. Open Systems Today, 164:38–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NSDb

[Ano94-148] Anonymous. New servers from Dell, ALR, and Compaq feature Windows NT support. Open Systems Today, 163:8–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NSDa

[Ano94-149] Anonymous. New single- and dual-processor RISC desktops introduced by FirePower and NeTpower. Open Systems Today, 163: 68–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NSO

[Ano94-150] Anonymous. New software opens OS/ 2 Notes access to data sources reached through Information Builders’ EDA/ SQL. Open Systems Today, 154:45–??, July 1994. ISSN 1061-0839.

Anonymous:1994:NTS

[Ano94-151] Anonymous. Northern Telecom sees emulation as a low-cost route to desktop interoperability. Open Systems Today, 158:78–??, September 1994. ISSN 1061-0839.

Anonymous:1994:NCA

[Ano94-152] Anonymous. Notes’ crossplatform abilities win Texaco’s management and users over to a single workgroup-computing platform. Open Systems Today, 163:84–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NHB

[Ano94-153] Anonymous. Novell held a briefing, God spoke, Mike came, and so did Mike, and Mike, and Mike, and. Open Systems Today, 163: 100–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NOP

[Ano94-154] Anonymous. Novell outlines plans to merge messaging software in the forthcoming NetWare 4.1. Open Systems Today, 165: 34–??, December 1994. ISSN 1061-0839.

Anonymous:1994:OAD

[Ano94-155] Anonymous. Object analysis and design capabilities take the spotlight at OOPSLA show. Open Systems Today, 163:58–??, November 1994. ISSN 1061-0839.

Anonymous:1994:OOS

[Ano94-156] Anonymous. Object-oriented operating system vendors are jousting for market primacy. For now, the lances consist mostly of words. Open Systems Today, 155: 42–??, August 1994. ISSN 1061-0839.

Anonymous:1994:OTG

[Ano94-157] Anonymous. Object-oriented tools: A guide to what’s available to application builders today. Open Systems Today, 155: 40–??, August 1994. ISSN 1061-0839.

Anonymous:1994:OMC

[Ano94-158] Anonymous. Objects and microkernels are cornerstones of modularity for data sharing in 32bit OSes. Open Systems Today, 163:64–??, November 1994. ISSN 1061-0839.

Anonymous:1994:OMG

[Ano94-159] Anonymous. Ontos middleware gives applications access to data via a C++ model. Open Systems Today, 163:14–??, November 1994. ISSN 1061-0839.

Anonymous:1994:OVW

[Ano94-160] Anonymous. OpenDoc versions for Windows and OS/ 2 have been shipped in alpha form by WordPerfect and IBM, respectively. Open Systems Today, 153: 10–??, July 1994. ISSN 1061-0839.

Anonymous:1994:OSW

[Ano94-161] Anonymous. OpenStep on Solaris will go into beta testing early next year. Open Systems Today, 165:5–??, December 1994. ISSN 1061-0839.

Anonymous:1994:OAI

[Ano94-162] Anonymous. Oracle and Avalon introduce integrated apps for financial and manufacturing use. Open Systems Today, 160:40–??, October 1994. ISSN 1061-0839.

Anonymous:1994:OAR

[Ano94-163] Anonymous. Oracle has announced remote management tools for databases and other products. Open Systems Today, 161:32–??, October 1994. ISSN 1061-0839.

Anonymous:1994:OIA

[Ano94-164] Anonymous. Oracle is announcing new wares this week at NetWorld-Interop, including a remote-access product and SQL middleware. Open Systems Today, 158:4–??, September 1994. ISSN 1061-0839.

Anonymous:1994:OPN

[Ano94-165] Anonymous. Ordering pizza over the net just had to happen. Open Systems Today, 157:78–??, September 1994. ISSN 1061-0839.

Anonymous:1994:OCH

[Ano94-166] Anonymous. OST compares how four database vendors would handle an application with complex data types. Among our questions: Which database would do the best job of voice recognition? Open Systems Today, 155:37–??, August 1994. ISSN 1061-0839.

Anonymous:1994:PDV

[Ano94-167] Anonymous. PC database vendors are bundling scripting languages and application development tools so users can customize local DBs easily. Open Systems Today, 159:22–??, September 1994. ISSN 1061-0839.

Anonymous:1994:PAG

[Ano94-168] Anonymous. PCExpo attendees groped for their checkbooks when Apple ran Windows-based applications on a Power Mac with a DOS- compatible card — but it was a demonstration only, and no product has been promised. Open Systems Today, 153:6–??, July 1994. ISSN 1061-0839.

Anonymous:1994:PDC

[Ano94-169] Anonymous. PCMCIA diskdrive cards get a capacity boost in products from MiniStor and Calluna. Open Systems Today, 155:46–??, August 1994. ISSN 1061-0839.

Anonymous:1994:PCM

[Ano94-170] Anonymous. Pharmaceutical company Marion Merrell Dow builds a self-managing database. Open Systems Today, 153:63–??, July 1994. ISSN 1061-0839.

Anonymous:1994:PTD

[Ano94-171] Anonymous. Platinum Technology’s data-staging system updates relational servers from a mainframe. Open Systems Today, 165:10–??, December 1994. ISSN 1061-0839.

Anonymous:1994:PMF

[Ano94-172] Anonymous. PowerPC-based machines flood the market, even though a hardware specification is not due until spring. Open Systems Today, 164:57–??, November 1994. ISSN 1061-0839.

Anonymous:1994:PRL

[Ano94-173] Anonymous. Product Review:

As a LAN OS, NetWare 4.1 will be hard to beat. Open Systems Today, 165:1–??, December 1994. ISSN 1061-0839.

Anonymous:1994:PRB

[Ano94-174] Anonymous. Product review:

Blyth Software’s Omnis 7 V.3 is a comprehensive development environment for building complex client-server GUI applications. Open Systems Today, 158: 72–??, September 1994. ISSN 1061-0839.

Anonymous:1994:PRI

[Ano94-175] Anonymous. Product review: integraph makes NT-NFS connections: PC users get access to NFS servers via PC-NFS for Windows NT, while DiskShare is an NFS server on the PC. Open Systems Today, 164:12–??, November 1994. ISSN 1061-0839.

Anonymous:1994:PRP

[Ano94-176] Anonymous. Product review: Pretty Good Privacy is a privacy advocate’s plan to head off the government’s bid to snoop on data. Open Systems Today, 162: 56–??, October 1994. ISSN 1061-0839.

Anonymous:1994:RRW

[Ano94-177] Anonymous. Role reversal: What happens when Microsoft stiffens up and IBM lets its hair down. Open Systems Today, 162: 82–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SIBb

[Ano94-178] Anonymous. SAS Institute’s business analysis software now supports several Unix platforms. Open Systems Today, 161:8–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SCS

[Ano94-179] Anonymous. Seven cellular service providers have founded the CDPD Forum to promote interoperability among Cellular Digital Packet Data products and services. Open Systems Today, 153:4–??, July 1994. ISSN 1061-0839.

Anonymous:1994:SEM

[Ano94-180] Anonymous. Several electronic mail front ends let users read and send E-mail from multiple vendors’ transports and servers. Open Systems Today, 154:28–??, July 1994. ISSN 1061-0839.

Anonymous:1994:SVW

[Ano94-181] Anonymous. Several vendors will introduce storage systems at next month’s Unix Expo. Open Systems Today, 157:54–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SGR

[Ano94-182] Anonymous. Silicon Graphics revs up the 3-D performance of its workstation and server lines. Open Systems Today, 155:5–??, August 1994. ISSN 1061-0839.

Anonymous:1994:SSI

[Ano94-183] Anonymous. Slow sales of its Improv spreadsheet app have prompted Lotus Development to discontinue the product. Open Systems Today, 157:34–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SWR

[Ano94-184] Anonymous. Speed’s the word as RISC processor vendors roll out chips — most with 64-bit designs — that are at least twice as fast as the Pentium. Open Systems Today, 160:84–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SIV

[Ano94-185] Anonymous. Spry has incorporated a version of Mosaic into its latest piece of Internet software. Open Systems Today, 157: 40–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SIBa

[Ano94-186] Anonymous. Spry’s Internet In A Box, a dial-up package for Net connectivity, is on the right track but isn’t as simple as it sounds. Open Systems Today, 161:41–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SUP

[Ano94-187] Anonymous. SQLWindows, Uniface and PowerBuilder clientserver tools make the most of object-oriented programming. Open Systems Today, 155:50–??, August 1994. ISSN 1061-0839.

Anonymous:1994:SPS

[Ano94-188] Anonymous. Start-up plans to ship multimedia networking products that eliminate net bottlenecks. Open Systems Today, 163:32–??, November 1994. ISSN 1061-0839.

Anonymous:1994:SOTe

[Ano94-189] Anonymous. Stop #10 on OST’s 11-part technical tour: Commerce on the Internet awaits the resolution of security and reliability issues. Open Systems Today, 162:38–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SCOb

[Ano94-190] Anonymous. Stop #11 concludes OST’s technical tour: Industry experts offer their views on networks of the future. Open Systems Today, 164:28–??, November 1994. ISSN 1061-0839.

Anonymous:1994:SOTa

[Ano94-191] Anonymous. Stop 5 on OSTs 11-part technical tour considers network routers — what they do, and whether they will be replaced by LAN switches. Open Systems Today, 153:28–??, July 1994. ISSN 1061-0839.

Anonymous:1994:SOTb

[Ano94-192] Anonymous. Stop 6 on OST’s 11part technical tour is the widearea network, where we look at low-cost ways to multiply effective bandwidth. Open Systems Today, 154:36–??, July 1994. ISSN 1061-0839.

Anonymous:1994:SOTc

[Ano94-193] Anonymous. Stop 7 on OST’s 11-part technical tour: A look at what happens when the Email message reaches its destinations. Open Systems Today, 157: 44–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SOTd

[Ano94-194] Anonymous. Stop #9 on OST’s 11-part technical tour: How network managers integrate SNA mainframes with distributed TCP/ IP nets. Open Systems Today, 160:52–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SPT

[Ano94-195] Anonymous. Sun plans today to unveil a turnkey server for quick and easy Internet access. Open Systems Today, 158:5–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SWWb

[Ano94-196] Anonymous. Sun’s Wabi Windows translation software matures, giving Unix users more flexibility. Open Systems Today, 162:64–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SUI

[Ano94-197] Anonymous. SunSoft upgrades its Interactive Unix operating system with NetWare connectivity, support for PCMCIA Ethernet cards and more Internet utilities. Open Systems Today, 154: 48–??, July 1994. ISSN 1061-0839.

Anonymous:1994:SSP

[Ano94-198] Anonymous. SunSoft’s Solar-Net PC-Admin software includes support for an emerging Internet standard, DHCP, which allows IP addresses to be dynamically allocated. Open Systems Today, 162:4–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SWWa

[Ano94-199] Anonymous. SunSoft’s Wabi 2.0 will support 24 applications, 10 more than before, but will need Microsoft Windows to do it. Open Systems Today, 161:35–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SGS

[Ano94-200] Anonymous. Support grows for SNA Server, as vendors introduce connectivity products for Microsoft’s mainframe-to-LAN software. Open Systems Today, 164:24–??, November 1994. ISSN 1061-0839.

Anonymous:1994:SWS

[Ano94-201] Anonymous. Sure, writing software is important, but writing code names can be just as vital. Open Systems Today, 153:74–??, July 1994. ISSN 1061-0839.

Anonymous:1994:SBM

[Ano94-202] Anonymous. Swiss Bank has made it a point to embrace leading-edge technology. Open Systems Today, 158:76–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SOE

[Ano94-203] Anonymous. Sybase offers an early look at an indexing product, due next summer, that will speed up queries against data warehouses. Open Systems Today, 165:50–??, December 1994. ISSN 1061-0839.

Anonymous:1994:SAT

[Ano94-204] Anonymous. Symantec adds TCP/ IP and Windows 95 support to remote-connectivity PC software. Open Systems Today, 163:42–??, November 1994. ISSN 1061-0839.

Anonymous:1994:TTW

[Ano94-205] Anonymous. Technology transforms a Washington law firm, as desktop applications, on-line databases, and electronic mail empower the barristers. Open Systems Today, 161:52–??, October 1994. ISSN 1061-0839.

Anonymous:1994:TIF

[Ano94-206] Anonymous. Tektronix introduces the first entry-level X terminals to have two PCMCIA slots. Open Systems Today, 158: 64–??, September 1994. ISSN 1061-0839.

Anonymous:1994:CIP

[Ano94-207] Anonymous. The console from IBM that provides a single interface for system and network management will soon enter beta tests. Open Systems Today, 160: 62–??, October 1994. ISSN 1061-0839.

Anonymous:1994:DUP

[Ano94-208] Anonymous. The DECpc Universal Platform now offers a choice among DEC’s three strategic operating systems — on entry-level hardware. Open Systems Today, 156:56–??, August 1994. ISSN 1061-0839.

Anonymous:1994:FTP

[Ano94-209] Anonymous. The first thirdparty product for IBM’s 25-Mbps ATM implementation is expected to be introduced today. Open Systems Today, 158:6–??, September 1994. ISSN 1061-0839.

Anonymous:1994:LSO

[Ano94-210] Anonymous. The Lisp and Smalltalk object-oriented languages are receiving integrated database support. Open Systems Today, 157:49–??, September 1994. ISSN 1061-0839.

Anonymous:1994:MIR

[Ano94-211] Anonymous. The market is rushing to switching technology, which threatens to eclipse hubs and routers. Open Systems Today, 160:44–??, October 1994. ISSN 1061-0839.

Anonymous:1994:NBG

[Ano94-212] Anonymous. The Netscape browser goes beyond standards, raising concerns over compatibility with future Web standards and products. Open Systems Today, 164:26–??, November 1994. ISSN 1061-0839.

Anonymous:1994:NRO

[Ano94-213] Anonymous. The new release of OS/ 2 is being positioned as a “here today” alternative to Windows 95. Open Systems Today, 161:5–??, October 1994. ISSN 1061-0839.

Anonymous:1994:OIG

[Ano94-214] Anonymous. The OpenDoc initiative gets a boost as major vendors unveil plans for tools — and the Object Management Group comes aboard. Open Systems Today, 160:5–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SCOa

[Ano94-215] Anonymous. The Santa Cruz Operation’s next Unix version will bring a mother lode of enhancements. Open Systems Today, 157:5–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SED

[Ano94-216] Anonymous. The Software Edge’s Defect Control System manages software bugs within a LAN environment or on a standalone PC. Open Systems Today, 157:58–??, September 1994. ISSN 1061-0839.

Anonymous:1994:SCW

[Ano94-217] Anonymous. The success of the chip will depend on a common hardware design, a freely licensed System 7, and native applications. Open Systems Today, 164:

41–??, November 1994. ISSN 1061-0839.

Anonymous:1994:SRE

[Ano94-218] Anonymous. To sell a refrigerator to an Eskimo, call it a wolfproof meat locker. Open Systems Today, 160:98–??, October 1994. ISSN 1061-0839.

Anonymous:1994:TPP

[Ano94-219] Anonymous. Two PPP packages for the Macintosh are now available: MacPPP and InterPPP. Open Systems Today, 159:46–??, September 1994. ISSN 1061-0839.

Anonymous:1994:USL

[Ano94-220] Anonymous. Unify says the latest release of its visual 4GL lets developers partition applications and distribute the parts across a network. Open Systems Today, 162:5–??, October 1994. ISSN 1061-0839.

Anonymous:1994:UBP

[Ano94-221] Anonymous. UnixWare branding program debuts, allowing other vendors to use the name. Open Systems Today, 160:6–??, October 1994. ISSN 1061-0839.

Anonymous:1994:ULT

[Ano94-222] Anonymous. Users are looking for tools that combine network management and system management functions; so far, no one product is ideal. Open Systems Today, 153:40–??, July 1994. ISSN 1061-0839.

Anonymous:1994:UTS

[Ano94-223] Anonymous. Users are taking a shine to recordable CDs, which are especially well suited to archiving. Open Systems Today, 156:52–??, August 1994. ISSN 1061-0839.

Anonymous:1994:UPB

[Ano94-224] Anonymous. Users have put buying decisions for Apple’s Power Macs on hold, because native software — especially office productivity apps — is not materializing as fast as promised. Open Systems Today, 154:8–??, July 1994. ISSN 1061-0839.

Anonymous:1994:VAI

[Ano94-225] Anonymous. Vendors aim to integrate network platforms, rolling out products that link systems managing different parts of a network. Open Systems Today, 164:56–??, November 1994. ISSN 1061-0839.

Anonymous:1994:VLC

[Ano94-226] Anonymous. Version 3.2 of Locus Computing’s Merge Windows-to-Unix software lets users cut and paste between the two environments. Open Systems Today, 156: 47–??, August 1994. ISSN 1061-0839.

Anonymous:1994:VTT

[Ano94-227] Anonymous. Version 5.0 of the Tuxedo transaction monitor supports more standards and platforms. Open Systems Today, 163:6–??, November 1994. ISSN 1061-0839.

Anonymous:1994:VGC

[Ano94-228] Anonymous. Visix’s Galaxy 2.0 cross-platform toolkit is sure to please developers of both multiple- and single-platform apps. Open Systems Today, 153: 50–??, July 1994. ISSN 1061-0839.

Anonymous:1994:VEC

[Ano94-229] Anonymous. Visual Edge’s Cross-Platform Toolst 1.0 takes a new approach to GUI portability, examined in OST’s in-depth review. Open Systems Today, 160: 70–??, October 1994. ISSN 1061-0839.

Anonymous:1994:WSW

[Ano94-230] Anonymous. Way out in the sticks, we like to talk about the weather, hydroelectric dams, and, of course, the latest in computing technology. Open Systems Today, 164:78–??, November 1994. ISSN 1061-0839.

Anonymous:1994:WSP

[Ano94-231] Anonymous. What’s in store for the paperless office: OST Labs looks at the impact of electronic forms packages. Open Systems Today, 154:58–??, July 1994. ISSN 1061-0839.

Anonymous:1994:WRD

[Ano94-232] Anonymous. While the RPCvs.-messaging debate rages, users should understand that it’s possible to combine the advantages of both. Open Systems Today, 156: 58–??, August 1994. ISSN 1061-0839.

Anonymous:1994:WFP

[Ano94-233] Anonymous. Why so few Pentium notebooks? Many vendors are waiting for a cooler version of the chip. Open Systems Today, 156:50–??, August 1994. ISSN 1061-0839.

Anonymous:1994:WMA

[Ano94-234] Anonymous. WordPerfect moved aggressively on the applicationsuite front with the announcement that users will be able to add third-party apps to the PerfectOffice 3.0 environment. Open Systems Today, 153:5–??, July 1994. ISSN 1061-0839.

Anonymous:1994:WHS

[Ano94-235] Anonymous. Workflow, a hard sell by itself, is finding success through integration in vertical apps. Open Systems Today, 164: 22–??, November 1994. ISSN 1061-0839.

Anonymous:1995:AMA

[Ano95a] Anonymous. 3Com has announced module additions to its LANPlex and SuperStack switching hubs, letting users have Token Ring switching for data center and workgroup environments. Open Systems Today, 168: 6–??, February 1995. ISSN 1061-0839.

Anonymous:1995:FSS

[Ano95b] Anonymous. A fundamental shift to switching is under way, an ASIC-based hardware clearly begins to push routing out of the LAN picture. Open Systems Today, 166:24–??, January 1995. ISSN 1061-0839.

Anonymous:1995:HPG

[Ano95c] Anonymous. A hardware package from Global Village offers no-frills connections for linking LANs to the Internet via a 28.8-Kbps modem or an ISDN line. Open Systems Today, 167:26–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NGF

[Ano95d] Anonymous. A new group formed by four software vendors will push for standards and user education about on-line analytical processing. Open Systems Today, 167:29–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NPW

[Ano95e] Anonymous. A new portable workstation from RDI challenges Sun’s Voyager with a lighter, more powerful notebook package. Open Systems Today, 168:36–??, February 1995. ISSN 1061-0839.

Anonymous:1995:NPL

[Ano95f] Anonymous. A new product from Legent will automate administrative operations across mainframe, Unix, and NetWare environments. Open Systems Today, 169:5–??, February 1995. ISSN 1061-0839.

Anonymous:1995:NRS

[Ano95g] Anonymous. A new RAID storage array from Encore Computer will let users share data between Unix systems and mainframes.

Open Systems Today, 169:6–??, February 1995. ISSN 1061-0839.

Anonymous:1995:NSP

[Ano95h] Anonymous. A new suite of programs from Oracle brings the Oracle7 database management system to PC and mobile laptop users. Open Systems Today, 167: 5–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NSW

[Ano95i] Anonymous. A new switch from Whitetree Network Technologies runs ATM and Ethernet through the same port. Open Systems Today, 169:22–??, February 1995. ISSN 1061-0839.

Anonymous:1995:NVS

[Ano95j] Anonymous. A new version of SunSoft’s SunPC emulation software includes the latest release of DOS and — for the first time — Windows. Open Systems Today, 168:4–??, February 1995. ISSN 1061-0839.

Anonymous:1995:WVC

[Ano95k] Anonymous. A Windows version of the Cornerstone visual dataanalysis tool was announced by BBN Software. Open Systems Today, 167:18–??, January 1995. ISSN 1061-0839.

Anonymous:1995:AFA

[Ano95l] Anonymous. Among a flurry of announcements of commercial World Wide Web servers are two that allow users to run Web sites and Windows NT. Open Systems Today, 168:5–??, February 1995. ISSN 1061-0839.

Anonymous:1995:EES

[Ano95m] Anonymous. An E-mail encryption standard should be in place by April, providing for the incorporation of Privacy Enhanced Mail into the MIME standard. Open Systems Today, 168:26–??, February 1995. ISSN 1061-0839.

Anonymous:1995:APP

[Ano95n] Anonymous. Apple pits Power Macs against Pentiums, adding memory cache and shipping them with faster processors and larger disk drives than previous models. Open Systems Today, 166:6–??, January 1995. ISSN 1061-0839.

Anonymous:1995:BPP

[Ano95o] Anonymous. Banyan plans to provide protocol adapters that will link its Unix-based Enterprise Network Services products to Oracle databases. Open Systems Today, 169:22–??, February 1995. ISSN 1061-0839.

Anonymous:1995:BVG

[Ano95p] Anonymous. Banyan Vines gets native TCP/ IP client support, plus an application package for ftp printer sharing and NFS support. Open Systems Today, 168: 22–??, February 1995. ISSN 1061-0839.

Anonymous:1995:BCB

[Ano95q] Anonymous. Blue Cross and Blue Shield of Oregon has turned to a modular architecture to integrate several legacy customerservice applications, while providing for easy addition of new modules as the system grows. Open Systems Today, 167:40–??, January 1995. ISSN 1061-0839.

Anonymous:1995:CIS

[Ano95r] Anonymous. Cabletron is set to roll out ATM switches that are designed to migrate users to “virtual” ATM networks that can be configured by software commands. Open Systems Today, 169:4–??, February 1995. ISSN 1061-0839.

Anonymous:1995:CBI

[Ano95s] Anonymous. Cadbury Beverages is building a system for project management based on Windows, with Unix servers, an Oracle database, and a Lotus Notes document management system melded into the solution. Open Systems Today, 166:44–??, January 1995. ISSN 1061-0839.

Anonymous:1995:CBL

[Ano95t] Anonymous. Canon brings out a lower-cost version of its Object.station Intel 486-based desktop for developing and running object-oriented applications under NextStep. Open Systems Today, 167:32–??, January 1995. ISSN 1061-0839.

Anonymous:1995:CSI

[Ano95u] Anonymous. Collabra Software is adding management tools to its Share group-conferencing application. Open Systems Today, 169:20–??, February 1995. ISSN 1061-0839.

Anonymous:1995:CRW

[Ano95v] Anonymous. Comparative review: Windows-based Smalltalks from Digitalk, IBM, and Parc-Place Systems go head-to-head. Open Systems Today, 169:34–??, February 1995. ISSN 1061-0839.

Anonymous:1995:EGT

[Ano95w] Anonymous. Electronics giant Texas Instruments tackles desktop resource management, basing its strategy on Microsoft’s SMS. Open Systems Today, 169:42–??, February 1995. ISSN 1061-0839.

Anonymous:1995:FCS

[Ano95x] Anonymous. Fault-resilient computing has a software side, which we investigate in the second of a two-part series on strategies to keep applications up and running. Open Systems Today, 166: 40–??, January 1995. ISSN 1061-0839.

Anonymous:1995:HWB

[Ano95y] Anonymous. HP-UX 10.0 will be unveiled this week, with newly tuned kernel and I/O paths, plus a multithreaded NFS implementation. Open Systems Today, 168:34–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IIN

[Ano95z] Anonymous. IBM introduces new versions of its job scheduling and trouble ticketing distributed tools. Open Systems Today, 169: 31–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IIW

[Ano95-27] Anonymous. IBM is working on a personal database that will run under OS/ 2 and Windows 95, and will let users access data on DB2 servers. Open Systems Today, 168:4–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IWB

[Ano95-28] Anonymous. IBM will build support for PeerLogic’s Pipes Platform communications middleware into IBM’s System Object Model. Open Systems Today, 169:4–??, February 1995. ISSN 1061-0839.

Anonymous:1995:INM

[Ano95-29] Anonymous. IBM’s new Model 3CT RS/ 6000 workstation provides the industry’s best floating-point price/performance numbers in the US$35,000 to US$50,000 price range. Open Systems Today, 169:32–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IVS

[Ano95-30] Anonymous. Ilog’s Views 2.0 supports interactive 2-D graphics and includes standard widgets. Open Systems Today, 169:28–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IRI

[Ano95-31] Anonymous. In-depth review:

Interleaf’s Cyberleaf converts word- processing documents to Web pages, mapping document styles to their HTML equivalents. Open Systems Today, 168:

38–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IRS

[Ano95-32] Anonymous. In-depth review: Sun’s Workshop 1.1 leads the field in development systems, thanks to features that are as good or better than the competition’s — and several that are unique. Open Systems Today, 167:34–??, January 1995. ISSN 1061-0839.

Anonymous:1995:IWa

[Ano95-33] Anonymous. Industry Watch. Open Systems Today, 166:11–??, January 1995. ISSN 1061-0839.

Anonymous:1995:IWc

[Ano95-34] Anonymous. Industry Watch. Open Systems Today, 167:12–??, January 1995. ISSN 1061-0839.

Anonymous:1995:IWd

[Ano95-35] Anonymous. Industry Watch. Open Systems Today, 168:12–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IWe

[Ano95-36] Anonymous. Industry Watch. Open Systems Today, 169:13–??, February 1995. ISSN 1061-0839.

Anonymous:1995:LIV

[Ano95-37] Anonymous. Live Internet voice communications via PCs is offered by three new software products. Open Systems Today, 169: 8–??, February 1995. ISSN 1061-

Anonymous:1995:LBA

[Ano95-38] Anonymous. Looking back at seven wild and wacky years of Open Systems Today and Unix Today! Open Systems Today, 169:58–??, February 1995. ISSN 1061-0839.

Anonymous:1995:Ma

[Ano95-39] Anonymous. Mail. Open Systems Today, 166:14–??, January 1995. ISSN 1061-0839.

Anonymous:1995:Mb

[Ano95-40] Anonymous. Mail. Open Systems Today, 167:14–??, January 1995. ISSN 1061-0839.

Anonymous:1995:Mc

[Ano95-41] Anonymous. Mail. Open Systems Today, 168:16–??, February 1995. ISSN 1061-0839.

Anonymous:1995:Md

[Ano95-42] Anonymous. Mail. Open Systems Today, 169:16–??, February 1995. ISSN 1061-0839.

Anonymous:1995:MPA

[Ano95-43] Anonymous. Mercury and ParcPlace announced new versions of Mercury’s automated testing tools optimized for ParcPlace’s VisualWorks Smalltalk development environment. Open Systems Today, 168:32–??, February 1995. ISSN 1061-0839.

Anonymous:1995:MIW

[Ano95-44] Anonymous. Microsoft is working toward a plug-and-play object-based architecture that offers easy-to-build custom business solutions. Open Systems Today, 166:17–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NBP

[Ano95-45] Anonymous. New book provides advertising advice for doing business on the Internet. Open Systems Today, 166:54–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NVT

[Ano95-46] Anonymous. New versions of two visual development systems let developers of PC-based GUIs use COBOL rather than 4GLs to define business logic. Open Systems Today, 167:4–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NAP

[Ano95-47] Anonymous. Novell has announced pricing for UnixWare 2.0, and the increases will now set the stage for a new round of competition in the Unix-on-Intel marketplace. Open Systems Today, 167:30–??, January 1995. ISSN 1061-0839.

Anonymous:1995:NWS

[Ano95-48] Anonymous. Novell will soon release a version of InForms that can be easily integrated with document management systems, databases, and E-mail. Open Systems Today, 166:20–??, January 1995. ISSN 1061-0839.

Anonymous:1995:OAP

[Ano95-49] Anonymous. On-line analytical processing toolsets provide new capabilities for analyzing complex data sets. Open Systems Today, 169:28–??, February 1995. ISSN 1061-0839.

Anonymous:1995:TUM

[Ano95-50] Anonymous. On Technology has upgraded its Meeting Maker XP scheduling app to run over the Internet with TCP/ IP and Winsock API support. Open Systems Today, 167:18–??, January 1995. ISSN 1061-0839.

Anonymous:1995:OSA

[Ano95-51] Anonymous. OpenLink Software has announced a product that allows the interchangeable development and deployment of multidatabase applications under Windows, Unix, and other OSecs. Open Systems Today, 166: 34–??, January 1995. ISSN 1061-0839.

Anonymous:1995:ORF

[Ano95-52] Anonymous. Oracle has released free tools for accessing its databases over the World Wide Web. Open Systems Today, 169:26–??, February 1995. ISSN 1061-0839.

Anonymous:1995:OIN

[Ano95-53] Anonymous. Oracle is now shipping encryption software for securing data on SQL*Net 2.1 networks — even data going to non-Oracle databases. Open Systems Today, 167:29–??, January 1995. ISSN 1061-0839.

Anonymous:1995:OSE

[Ano95-54] Anonymous. Our senior editor takes a look back on the past five years of open systems. Open Systems Today, 168:54–??, February 1995. ISSN 1061-0839.

Anonymous:1995:PIS

[Ano95-55] Anonymous. Pacer is set to announce server-based middleware that will let Windows and Mac apps access multiple vendors’ databases. Open Systems Today, 168:8–??, February 1995. ISSN 1061-0839.

Anonymous:1995:PAB

[Ano95-56] Anonymous. Porting apps between Windows and Unix is easier with packages that are specifically designed for that purpose. Open Systems Today, 166:32–??, January 1995. ISSN 1061-0839.

Anonymous:1995:PRI

[Ano95-57] Anonymous. Product review: IXI’s Eye2eye links. Open Systems Today, 166:38–??, January 1995. ISSN 1061-0839.

Anonymous:1995:SIT

[Ano95-58] Anonymous. Secure Internet transmissions are featured in Netscape Communications’ new World Wide Web server software. Open Systems Today, 166:6–??, January 1995. ISSN 1061-0839.

Anonymous:1995:SVW

[Ano95-59] Anonymous. Several vendors are working on software that will let users run World Wide Web pages on Windows servers. Open Systems Today, 166:30–??, January 1995. ISSN 1061-0839.

Anonymous:1995:SES

[Ano95-60] Anonymous. Software from Enterprise Systems Management will let administrator groups collectively manage multiple name services. Open Systems Today, 169:31–??, February 1995. ISSN 1061-0839.

Anonymous:1995:STF

[Ano95-61] Anonymous. STL Toolkit, the first commercial implementation of ANSI- standard software that eases the writing of portable C++ code, is now available. Open Systems Today, 167:6–??, January 1995. ISSN 1061-0839.

Anonymous:1995:TII

[Ano95-62] Anonymous. TCP/ IP is gaining a big following among PC LAN users, who are looking to more closely integrate all parts of the enterprise. Open Systems Today, 167:20–??, January 1995. ISSN 1061-0839.

Anonymous:1995:TSL

[Ano95-63] Anonymous. Third-party support for Lotus Notes takes off as the user following grows; IBM, Apple, and AT and T all announce bundling or integration plans. Open Systems Today, 168: 20–??, February 1995. ISSN 1061-0839.

Anonymous:1995:QTR

[Ano95-64] Anonymous. This quarter’s TPC results show little change, as vendors sit out a few rounds while benchmarks get an overhaul, and “benchmark special” disappear.

Open Systems Today, 166:37–??, January 1995. ISSN 1061-0839.

Anonymous:1995:TPA

[Ano95-65] Anonymous. Trinzic plans to announce a tool that lets Notes users access a wide variety of relational databases from within the Notes environment. Open Systems Today, 167:4–??, January 1995. ISSN 1061-0839.

Anonymous:1995:VOD

[Ano95-66] Anonymous. Version 3.5 of Objectivity/DB lets objects created with SQL, C++, and ParcPlace Smalltalk interoperate within an application. Open Systems Today, 167:8–??, January 1995. ISSN 1061-0839.

Anonymous:1995:WPI

[Ano95-67] Anonymous. Wacky playtime is the right kind of break for brainweary technical conference-goers. Open Systems Today, 167:54–??, January 1995. ISSN 1061-0839.

Anonymous:1995:WCS

[Ano95-68] Anonymous. When a comet slammed into Jupiter last year, the Space Telescope Science Institute began receiving tons of data from the watching eye of the Hubble telescope. Open Systems Today, 168:42–??, February 1995. ISSN 1061-0839.

Anonymous:1995:WIS

[Ano95-69] Anonymous. Windows 95 and increased support for Windows clients will make it easier to manage PCs as network clients. Open Systems Today, 166:36–??, January 1995. ISSN 1061-0839.

Atkins:1994:UOb

[Atk94a] Sally Atkins. THe User Odyssey.

Open Systems Today, 155:14–??, August 1994. ISSN 1061-0839.

Atkins:1994:UOg

[Atk94b] Sally Atkins. The User Odyssey.

Open Systems Today, 160:84–??, October 1994. ISSN 1061-0839.

Atkins:1994:UOi

[Atk94c] Sally Atkins. The User Odyssey.

Open Systems Today, 162:68–??, October 1994. ISSN 1061-0839.

Atkins:1994:UOj

[Atk94d] Sally Atkins. The User Odyssey.

Open Systems Today, 163:86–??, November 1994. ISSN 1061-0839.

Atkins:1994:UOk

[Atk94e] Sally Atkins. The User Odyssey.

Open Systems Today, 164:65–??, November 1994. ISSN 1061-0839.

Atkins:1994:UOl

[Atk94f] Sally Atkins. The User Odyssey.

Open Systems Today, 165:74–??, December 1994. ISSN 1061-0839.

Atkins:1994:UOa

[Atk94g] Sally Atkins. User Odyssey.

Open Systems Today, 154:73–??, July 1994. ISSN 1061-0839.

Atkins:1994:UOc

[Atk94h] Sally Atkins. User Odyssey.

Open Systems Today, 156:68–??, August 1994. ISSN 1061-0839.

Atkins:1994:UOd

[Atk94i] Sally Atkins. User Odyssey. Open Systems Today, 157:74–??, September 1994. ISSN 1061-0839.

Atkins:1994:UOe

[Atk94j] Sally Atkins. User Odyssey. Open Systems Today, 158:78–??, September 1994. ISSN 1061-0839.

Atkins:1994:UOf

[Atk94k] Sally Atkins. User Odyssey. Open Systems Today, 159:56–??, September 1994. ISSN 1061-0839.

Atkins:1994:UOh

[Atk94l] Sally Atkins. User Odyssey. Open Systems Today, 161:53–??, October 1994. ISSN 1061-0839.

Atkins:1995:UOa

[Atk95a] Sally Atkins. The User Odyssey. Open Systems Today, 166:50–??, January 1995. ISSN 1061-0839.

Atkins:1995:UOb

[Atk95b] Sally Atkins. The User Odyssey. Open Systems Today, 168:43–??, February 1995. ISSN 1061-0839.

Atkins:1995:UOc

[Atk95c] Sally Atkins. The User Odyssey. Open Systems Today, 169:43–??, February 1995. ISSN 1061-0839.

Azzara:1994:MVj

[Azz94a] Mike Azzara. Mike’s View. Open Systems Today, 162:26–??, October 1994. ISSN 1061-0839.

Azzara:1994:MVk

[Azz94b] Mike Azzara. Mike’s View. Open Systems Today, 163:28–??, November 1994. ISSN 1061-0839.

Azzara:1994:MVl

[Azz94c] Mike Azzara. Mike’s View. Open Systems Today, 164:16–??, November 1994. ISSN 1061-0839.

Azzara:1994:MVm

[Azz94d] Mike Azzara. Mike’s View. Open Systems Today, 165:28–??, December 1994. ISSN 1061-0839.

Azzara:1994:MVa

[Azz94e] Mike Azzara. My View. Open Systems Today, 153:20–??, July 1994. ISSN 1061-0839.

Azzara:1994:MVb

[Azz94f] Mike Azzara. My View. Open Systems Today, 154:26–??, July 1994. ISSN 1061-0839.

Azzara:1994:MVc

[Azz94g] Mike Azzara. My View. Open Systems Today, 155:26–??, August 1994. ISSN 1061-0839.

Azzara:1994:MVd

[Azz94h] Mike Azzara. My View. Open Systems Today, 156:24–??, August 1994. ISSN 1061-0839.

Azzara:1994:MVe

[Azz94i] Mike Azzara. My View. Open Systems Today, 157:26–??, September 1994. ISSN 1061-0839.

Azzara:1994:MVf

[Azz94j] Mike Azzara. My View. Open Systems Today, 158:32–??, September 1994. ISSN 1061-0839.

Azzara:1994:MVg

[Azz94k] Mike Azzara. My View. Open Systems Today, 159:20–??, September 1994. ISSN 1061-0839.

Azzara:1994:MVh

[Azz94l] Mike Azzara. My View. Open Systems Today, 160:32–??, October 1994. ISSN 1061-0839.

Azzara:1994:MVi

[Azz94m] Mike Azzara. My View. Open Systems Today, 161:21–??, October 1994. ISSN 1061-0839.

Azzara:1995:MVa

[Azz95a] Mike Azzara. Mike’s View. Open Systems Today, 166:14–??, January 1995. ISSN 1061-0839.

Azzara:1995:MVb

[Azz95b] Mike Azzara. Mike’s View. Open Systems Today, 167:14–??, January 1995. ISSN 1061-0839.

Azzara:1995:MVc

[Azz95c] Mike Azzara. Mike’s View. Open Systems Today, 168:16–??, February 1995. ISSN 1061-0839.

Azzara:1995:MVd

[Azz95d] Mike Azzara. Mike’s View. Open Systems Today, 169:16–??, February 1995. ISSN 1061-0839.

Bean:1994:Ea

[Bea94a] Gary Bean. Explorations. Open Systems Today, 154:70–??, July 1994. ISSN 1061-0839.

Bean:1994:Eb

[Bea94b] Gary Bean. Explorations. Open Systems Today, 156:64–??, August 1994. ISSN 1061-0839.

Bean:1994:Ec

[Bea94c] Gary Bean. Explorations. Open Systems Today, 157:63–??, September 1994. ISSN 1061-0839.

Bean:1994:Ed

[Bea94d] Gary Bean. Explorations. Open Systems Today, 158:74–??, September 1994. ISSN 1061-0839.

Bean:1994:Ee

[Bea94e] Gary Bean. Explorations. Open Systems Today, 159:50–??, September 1994. ISSN 1061-0839.

Bean:1994:Ef

[Bea94f] Gary Bean. Explorations. Open Systems Today, 160:79–??, October 1994. ISSN 1061-0839.

Bean:1994:Eg

[Bea94g] Gary Bean. Explorations. Open Systems Today, 161:44–??, October 1994. ISSN 1061-0839.

Bean:1994:Eh

[Bea94h] Gary Bean. Explorations. Open Systems Today, 162:57–??, October 1994. ISSN 1061-0839.

Bean:1994:Ei

[Bea94i] Gary Bean. Explorations. Open Systems Today, 163:82–??, November 1994. ISSN 1061-0839.

Bean:1994:Ej

[Bea94j] Gary Bean. Explorations. Open Systems Today, 164:62–??, November 1994. ISSN 1061-0839.

Bean:1994:Ek

[Bea94k] Gary Bean. Explorations. Open Systems Today, 165:66–??, December 1994. ISSN 1061-0839.

Bean:1995:Ea

[Bea95a] Gary Bean. Explorations. Open Systems Today, 166:43–??, January 1995. ISSN 1061-0839.

Bean:1995:Eb

[Bea95b] Gary Bean. Explorations. Open Systems Today, 168:41–??, February 1995. ISSN 1061-0839.

Bean:1995:Ec

[Bea95c] Gary Bean. Explorations. Open Systems Today, 169:41–??, February 1995. ISSN 1061-0839.

Feibus:1994:SPa

[Fei94a] Andy Feibus. Standards in Progress. Open Systems Today, 153:16–??, July 1994. ISSN 1061-0839.

Feibus:1994:SPb

[Fei94b] Andy Feibus. Standards In Progress. Open Systems Today, 154:20–??, July 1994. ISSN 1061-0839.

Feibus:1994:SPc

[Fei94c] Andy Feibus. Standards in Progress. Open Systems Today, 155:20–??, August 1994. ISSN 1061-0839.

Feibus:1994:SPe

[Fei94d] Andy Feibus. Standards In Progress. Open Systems Today, 157:20–??, September 1994. ISSN 1061-0839.

Feibus:1994:SPf

[Fei94e] Andy Feibus. Standards in Progress. Open Systems Today, 158:26–??, September 1994. ISSN 1061-0839.

Feibus:1994:SPg

[Fei94f] Andy Feibus. Standards in Progress. Open Systems Today, 159:14–??, September 1994. ISSN 1061-0839.

Feibus:1994:SPh

[Fei94g] Andy Feibus. Standards in Progress. Open Systems Today, 160:26–??, October 1994. ISSN 1061-0839.

Feibus:1994:SPi

[Fei94h] Andy Feibus. Standards in Progress. Open Systems Today, 161:18–??, October 1994. ISSN 1061-0839.

Feibus:1994:SPj

[Fei94i] Andy Feibus. Standards in Progress. Open Systems Today, 162:20–??, October 1994. ISSN 1061-0839.

Feibus:1994:SPk

[Fei94j] Andy Feibus. Standards In Progress. Open Systems Today, 163:20–??, November 1994. ISSN 1061-0839.

Feibus:1994:SPl

[Fei94k] Andy Feibus. Standards In Progress. Open Systems Today, 164:12–??, November 1994. ISSN 1061-0839.

Feibus:1994:SPm

[Fei94l] Andy Feibus. Standards in Progress. Open Systems Today, 165:20–??, December 1994. ISSN 1061-0839.

Feibus:1994:SPd

[Fei94m]  Andy Feibus.  Stantards in

Progress. Open Systems Today, 156:20–??, August 1994. ISSN 1061-0839.

Feibus:1995:SPa

[Fei95a] Andy Feibus. Standards In Progress. Open Systems Today, 166:10–??, January 1995. ISSN 1061-0839.

Feibus:1995:SPb

[Fei95b] Andy Feibus. Standards In Progress. Open Systems Today, 167:10–??, January 1995. ISSN 1061-0839.

Feibus:1995:SPc

[Fei95c] Andy Feibus. Standards In Progress. Open Systems Today, 168:10–??, February 1995. ISSN 1061-0839.

Feibus:1995:SPd

[Fei95d] Andy Feibus. Standards In Progress. Open Systems Today, 169:14–??, February 1995. ISSN 1061-0839.

Gaskin:1994:Na

[Gas94a] James Gaskin. Networking. Open Systems Today, 153:26–??, July 1994. ISSN 1061-0839.

Gaskin:1994:Nb

[Gas94b] James Gaskin. Networking. Open Systems Today, 154:32–??, July 1994. ISSN 1061-0839.

Gaskin:1994:Nc

[Gas94c] James Gaskin. Networking. Open Systems Today, 155:32–??, August 1994. ISSN 1061-0839.

Gaskin:1994:Nd

[Gas94d] James Gaskin. Networking. Open Systems Today, 156:34–??, August 1994. ISSN 1061-0839.

Gaskin:1994:Ne

[Gas94e] James Gaskin. Networking. Open Systems Today, 157:38–??, September 1994. ISSN 1061-0839.

Gaskin:1994:Nf

[Gas94f] James Gaskin. Networking. Open Systems Today, 158:40–??, September 1994. ISSN 1061-0839.

Gaskin:1994:Ng

[Gas94g] James Gaskin. Networking. Open Systems Today, 159:30–??, September 1994. ISSN 1061-0839.

Gaskin:1994:Nh

[Gas94h] James Gaskin. Networking. Open Systems Today, 160:44–??, October 1994. ISSN 1061-0839.

Gaskin:1994:Ni

[Gas94i] James Gaskin. Networking. Open Systems Today, 161:30–??, October 1994. ISSN 1061-0839.

Gaskin:1994:Nj

[Gas94j] James Gaskin. Networking. Open Systems Today, 162:37–??, October 1994. ISSN 1061-0839.

Gaskin:1994:Nk

[Gas94k] James Gaskin. Networking. Open Systems Today, 163:46–??, November 1994. ISSN 1061-0839.

Gaskin:1994:Nl

[Gas94l] James Gaskin. Networking. Open Systems Today, 164:26–??, November 1994. ISSN 1061-0839.

Gaskin:1994:Nm

[Gas94m] James Gaskin. Networking. Open Systems Today, 165:42–??, December 1994. ISSN 1061-0839.

Gaskin:1995:Na

[Gas95a] James Gaskin. Networking. Open Systems Today, 166:29–??, January 1995. ISSN 1061-0839.

Gaskin:1995:Nb

[Gas95b] James Gaskin. Networking. Open Systems Today, 167:24–??, January 1995. ISSN 1061-0839.

Gaskin:1995:Nc

[Gas95c] James Gaskin. Networking. Open Systems Today, 168:24–??, February 1995. ISSN 1061-0839.

Gaskin:1995:Nd

[Gas95d] James Gaskin. Networking. Open Systems Today, 169:24–??, February 1995. ISSN 1061-0839.

Hunter:1994:SAa

[Hun94a] Bruce Hunter. Systems Administration. Open Systems Today, 153:56–??, July 1994. ISSN 1061-0839.

Hunter:1994:SAb

[Hun94b] Bruce Hunter. Systems Administration. Open Systems Today, 154:64–??, July 1994. ISSN 1061-0839.

Hunter:1994:SAc

[Hun94c] Bruce Hunter. Systems Administration. Open Systems Today, 156:62–??, August 1994. ISSN 1061-0839.

Hunter:1994:SAd

[Hun94d] Bruce Hunter. Systems Administration. Open Systems Today, 157:59–??, September 1994. ISSN 1061-0839.

Hunter:1994:SAe

[Hun94e] Bruce Hunter. Systems Administration. Open Systems Today, 158:75–??, September 1994. ISSN 1061-0839.

Hunter:1994:SAf

[Hun94f] Bruce Hunter. Systems Administration. Open Systems Today, 159:47–??, September 1994. ISSN 1061-0839.

Hunter:1994:SAg

[Hun94g] Bruce Hunter. Systems Administration. Open Systems Today, 160:74–??, October 1994. ISSN 1061-0839.

Hunter:1994:SAh

[Hun94h] Bruce Hunter. Systems Administration. Open Systems Today, 161:51–??, October 1994. ISSN 1061-0839.

Hunter:1994:SAi

[Hun94i] Bruce Hunter. Systems Administration. Open Systems Today, 162:60–??, October 1994. ISSN 1061-0839.

Hunter:1994:SAj

[Hun94j] Bruce Hunter. Systems Administration. Open Systems Today, 163:74–??, November 1994. ISSN 1061-0839.

Hunter:1994:SAk

[Hun94k] Bruce Hunter. Systems Administration. Open Systems Today, 164:60–??, November 1994. ISSN 1061-0839.

Hunter:1994:SAl

[Hun94l] Bruce Hunter. Systems Administration. Open Systems Today, 165:63–??, December 1994. ISSN 1061-0839.

Hunter:1995:SAa

[Hun95a] Bruce Hunter. Systems Administration. Open Systems Today, 166:42–??, January 1995. ISSN 1061-0839.

Hunter:1995:SAb

[Hun95b] Bruce Hunter. Systems Administration. Open Systems Today, 167:36–??, January 1995. ISSN 1061-0839.

Hunter:1995:SAc

[Hun95c] Bruce Hunter. Systems Administration. Open Systems Today, 168:40–??, February 1995. ISSN 1061-0839.

Hunter:1995:SAd

[Hun95d] Bruce Hunter. Systems Administration. Open Systems Today, 169:39–??, February 1995. ISSN 1061-0839.

Irsfeld:1994:MVa

[Irs94a] Mitch Irsfeld. My View. Open Systems Today, 162:26–??, October 1994. ISSN 1061-0839.

Irsfeld:1994:MVb

[Irs94b] Mitch Irsfeld. My View. Open Systems Today, 163:28–??, November 1994. ISSN 1061-0839.

Irsfeld:1994:MVc

[Irs94c] Mitch Irsfeld. My View. Open Systems Today, 164:16–??, November 1994. ISSN 1061-0839.

Irsfeld:1994:MVd

[Irs94d] Mitch Irsfeld. My View. Open Systems Today, 165:28–??, December 1994. ISSN 1061-0839.

Irsfeld:1995:MVa

[Irs95a] Mitch Irsfeld. My View. Open Systems Today, 166:14–??, January 1995. ISSN 1061-0839.

Irsfeld:1995:MVb

[Irs95b] Mitch Irsfeld. My View. Open Systems Today, 167:14–??, January 1995. ISSN 1061-0839.

Irsfeld:1995:MVc

[Irs95c] Mitch Irsfeld. My View. Open Systems Today, 168:16–??, February 1995. ISSN 1061-0839.

Irsfeld:1995:MVd

[Irs95d] Mitch Irsfeld. My View. Open Systems Today, 169:16–??, February 1995. ISSN 1061-0839.

Levit:1994:SR

[Lev94a] Jason Levit. Shrink Rap. Open Systems Today, 153:54–??, July 1994. ISSN 1061-0839.

Levitt:1994:SRa

[Lev94b] Jason Levitt. Shrink Rap. Open Systems Today, 154:68–??, July 1994. ISSN 1061-0839.

Levitt:1994:SRb

[Lev94c] Jason Levitt. Shrink Rap. Open Systems Today, 156:60–??, August 1994. ISSN 1061-0839.

Levitt:1994:SRc

[Lev94d] Jason Levitt. Shrink Rap. Open Systems Today, 157:62–??, September 1994. ISSN 1061-0839.

Levitt:1994:SRd

[Lev94e] Jason Levitt. Shrink Rap. Open Systems Today, 159:53–??, September 1994. ISSN 1061-0839.

Levitt:1994:SRe

[Lev94f] Jason Levitt. Shrink Rap. Open Systems Today, 160:78–??, October 1994. ISSN 1061-0839.

Levitt:1994:SRf

[Lev94g] Jason Levitt. Shrink Rap. Open Systems Today, 161:48–??, October 1994. ISSN 1061-0839.

Levitt:1994:SRg

[Lev94h] Jason Levitt. Shrink Rap. Open Systems Today, 162:62–??, October 1994. ISSN 1061-0839.

Levitt:1994:SRh

[Lev94i] Jason Levitt. Shrink Rap. Open Systems Today, 163:81–??, November 1994. ISSN 1061-0839.

Levitt:1994:SRi

[Lev94j] Jason Levitt. Shrink Rap. Open Systems Today, 164:63–??, November 1994. ISSN 1061-0839.

Levitt:1994:SRj

[Lev94k] Jason Levitt. Shrink Rap. Open Systems Today, 165:70–??, December 1994. ISSN 1061-0839.

Levitt:1995:SRa

[Lev95a] Jason Levitt. Shrink Rap. Open Systems Today, 166:39–??, January 1995. ISSN 1061-0839.

Levitt:1995:SRb

[Lev95b] Jason Levitt. Shrink Rap. Open Systems Today, 167:38–??, January 1995. ISSN 1061-0839.

Levitt:1995:SRc

[Lev95c] Jason Levitt. Shrink Rap. Open Systems Today, 168:39–??, February 1995. ISSN 1061-0839.

Levitt:1995:SRd

[Lev95d] Jason Levitt. Shrink Rap. Open Systems Today, 169:40–??, February 1995. ISSN 1061-0839.

Lytton:1994:OSAa

[Lyt94a] Nina Lytton. Open Systems Advisor. Open Systems Today, 153: 14–??, July 1994. ISSN 1061-0839.

Lytton:1994:OSAb

[Lyt94b] Nina Lytton. Open Systems Advisor. Open Systems Today, 154: 16–??, July 1994. ISSN 1061-0839.

Lytton:1994:OSAc

[Lyt94c] Nina Lytton. Open Systems Advisor. Open Systems Today, 155: 17–??, August 1994. ISSN 1061-0839.

Lytton:1994:OSAd

[Lyt94d] Nina Lytton. Open Systems Advisor. Open Systems Today, 156: 16–??, August 1994. ISSN 1061-0839.

Lytton:1994:OSAe

[Lyt94e] Nina Lytton. Open Systems Advisor. Open Systems Today, 157: 16–??, September 1994. ISSN 1061-0839.

Lytton:1994:OSAf

[Lyt94f] Nina Lytton. Open Systems Advisor. Open Systems Today, 158: 24–??, September 1994. ISSN 1061-0839.

Lytton:1994:OSAg

[Lyt94g] Nina Lytton. Open Systems Advisor. Open Systems Today, 159: 12–??, September 1994. ISSN 1061-0839.

Lytton:1994:OSAh

[Lyt94h] Nina Lytton. Open Systems Advisor. Open Systems Today, 160: 22–??, October 1994. ISSN 1061-0839.

Lytton:1994:OSAi

[Lyt94i] Nina Lytton. Open Systems Advisor. Open Systems Today, 161: 16–??, October 1994. ISSN 1061-0839.

Lytton:1994:OSAj

[Lyt94j] Nina Lytton. Open Systems Advisor. Open Systems Today, 162: 16–??, October 1994. ISSN 1061-0839.

Lytton:1994:OSAk

[Lyt94k] Nina Lytton. Open Systems Advisor. Open Systems Today, 163: 17–??, November 1994. ISSN 1061-0839.

Lytton:1994:OSAl

[Lyt94l] Nina Lytton. Open Systems Advisor. Open Systems Today, 164: 10–??, November 1994. ISSN 1061-0839.

Lytton:1994:OSAm

[Lyt94m] Nina Lytton. Open Systems Advisor. Open Systems Today, 165: 16–??, December 1994. ISSN 1061-0839.